### v1

# BeFA: A General Behavior-driven Feature Adapter for Multimedia Recommendation

Abstract.
Multimedia recommender systems focus on utilizing behavioral information and content information to model user preferences. Typically, it employs pre-trained feature encoders to extract content features, then fuses them with behavioral features. However, pre-trained feature encoders often extract features from the entire content simultaneously, including excessive preference-irrelevant details. We speculate that it may result in the extracted features not containing sufficient features to accurately reflect user preferences.
To verify our hypothesis, we introduce an attribution analysis method for visually and intuitively analyzing the content features. The results indicate that certain products’ content features exhibit the issues of information drift and information omission, reducing the expressive ability of features. Building upon this finding, we propose an effective and efficient general Behavior-driven Feature Adapter (BeFA) to tackle these issues. This adapter reconstructs the content feature with the guidance of behavioral information, enabling content features accurately reflecting user preferences. Extensive experiments demonstrate the effectiveness of the adapter across all multimedia recommendation methods. The code will be publicly available upon the paper’s acceptance.
Multimodal Recommendation, Feature Adapter, Information Drift, Information Omission
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06submissionid: 123-A56-BU3
Figure 1. Illustration of content features that do not accurately reflect the users’ preferences. Excessive irrelevant information hinders the recommender system’s ability to effectively model the users’ true preferences.
1. Introduction
Recommender systems have gained widespread adoption across various domains, aiming to assist users in discovering information that aligns with their preferences (vldb1, ; vldb2, ). In the case of multimedia platforms, the abundance of data resources provides recommender systems with increased opportunities to accurately model user preferences (survey, ).
Existing multimedia recommendation methods typically involves two main steps (w2go, ; survey, ). First, a pre-trained feature encoder is employed to capture content features from diverse modalities. Subsequently, these content features are fused with behavioral features to obtain user preference representations. Recently, researchers have focused on enhancing the quality of these representations through self-supervised learning. For instance, SLMRec (slmrec, ) investigates potential relationships between modalities through the use of contrastive learning, thereby obtaining a powerful representation. BM3 (bm3, ) utilizes a dropout strategy to construct multiple views and reconstructs the interaction graph, incorporating intra- and inter-modality contrastive loss to facilitate effective representation learning. MICRO (micro, ) and MGCN (mgcn, ) maximize the mutual information between content features and behavioral features with a self-supervised auxiliary task, and have achieved excellent performance.
Figure 2. Results of visualisation attribution analysis on the TMALL dataset.(a) is the original product image and (b) is its corresponding heatmap.
Despite the good progress made by existing methods in utilizing content and behavioral information more effectively, a crucial yet easily overlooked problem arises: Are content features obtained from pre-trained encoders containing sufficient features to reflect user preferences? Intuitively, multimedia content inherently exhibits the characteristic of low informational value density, where a significant portion of the presented information may be irrelevant to the users’ focus (micro, ; survey, ). Pre-trained feature encoders extract information from the entire content simultaneously, which can result in content features that do not truly reflect the users’ preferences (as shown in Figure 1). Fusing these irrelevant content features with behavioral features may mislead user preference modeling, resulting in suboptimal recommendation performance.
To answer this question, we introduce a similarity-based attribution analysis method for visualizing and intuitively analyzing the content features. This method evaluates the extent to which content features can reflect user preferences, enabling researchers to visually assess the quality of content features for the first time. The results indicate that not all products’ content features accurately reflect user preferences. Due to the presence of irrelevant information, certain products’ content features exhibit the issues of information drift and information omission. As shown in Figure 2, some products’ content features do not include information about the products that users are interested in, but instead erroneously include information about unrelated items, a phenomenon we term information drift. There are also some products’ content features that omit certain key details of the products, which we refer to as information omission. These issues ultimately prevent recommender systems from accurately modeling user preferences. To verify the rationality of these findings, we provide a theoretical analysis and explanation. Furthermore, we propose a plug-and-play general Behavior-driven Feature Adapter (BeFA) to address the discovered issues. This adapter effectively decouples, filters and reconstructs content features, leveraging behavioral information as a guide to obtain more precise representations of content information. Extensive experiments demonstrate the adapter’s effectiveness across various multimedia recommendation methods and feature encoders.
Our main contributions can be summarized as follows:
  • We introduce a similarity-based visual attribution method, which enables researchers to visually analyze the quality of content features for the first time.
  • We experimentally revealed the issues of information drift and information omission in content features. We also provide a theoretical analysis to validate the rationality of these findings.
  • We propose a general behavior-driven feature adapter, which obtain more precise content representation through decoupling and reconstructing content features.
Figure 3. Pipeline of the proposed attribution analysis method.
2. RELATED works
2.1. Multimedia Recommendation
Collaborative filtering (CF)-based approaches have achieved great success in recommender systems, relying on behavioral similarities to make top-k recommendations (cf1, ; cf2, ). Since user preferences are usually influenced by multimodal information, it facilitates researchers to integrate multimodal content information into CF-based approaches with the aim of improvement. Typically, content features are first extracted using pre-trained neural networks and then fused with behavioral features to enhance preference modeling (vbpr, ; w2go, ). Considering that behavioral information can be naturally modeled as a bipartite graph structure, researchers begin work on recommender systems based on Graph Convolutional Network(GCN) (gcn1, ; gcn2, ). For instance, MMGCN citemmgcn constructs modality-specific user-item interaction graphs to model user preferences specific to each modality. LATTICE (lattice, ) adds links between items with similar modality characteristics and create a separate item-item graph for each modality. FREEDOM (FREEDOM, ) constructs item-item graphs to help learn item semantic relations and reduces the graphs according to the sensitivity edge pruning technique. Meanwhile, researchers began to focus on enhancing the quality of representations through self-supervised learning. BM3 (bm3, ) learns user and item representations by reconstructing the user-item interaction graph and aligning modality features from both inter and intra-modality perspectives. MGCN (mgcn, ) comprehensively models user preferences by adaptively learning the relative importance of different modality features.
However, existing methods usually directly utilize features extracted from pre-trained feature encoders as side information for each item. The pre-trained feature encoders are trained on large-scale datasets collected from the natural world. The features of natural data and e-commerce data are very different. In e-commerce scenarios, the content information often contains a lot of irrelevant information and exhibits the characteristic of low value density. For example, only a small portion of the content image contain relevant information, while the rest is considered noise. This disparity leads to a substantial amount of noise in the contetn features extracted by pre-trained encoders when applied to multimodal recommender systems. Consequently, the content features do not contain enough features to sufficiently reflect users’ preferences. Fusing these irrelevant content features with behavioral features may mislead user preference modeling, resulting in suboptimal recommendation performance.
2.2. Attribution Analysis
Methods for modality feature attribution analysis have been widely studied, aiding in the visualization of modality feature content. Class Activation Mapping (CAM) (CAM, ) is a method used to visualize the decision-making process of convolutional neural networks by generating heatmaps that highlight regions of an image contributing to a particular class prediction. Grad-CAM (gradcam, ) enhances this by using the gradient of the target class with respect to the feature maps of a specific convolutional layer, creating heatmaps that can be applied to various network architectures without altering the network structure. Score-CAM (scoreCAM, ) further improves upon Grad-CAM by weighting the activation map using output scores instead of gradient information, thus avoiding gradient-related problems. In multimodal recommender systems, t-SNE (tsne, ) is often used to visualize the distribution of samples in high-dimensional feature spaces, providing insights into the relationships and distributions among different modality features.
However, current methods (cvpr, ; gScoreCAM, ) for analyzing image feature weights predominantly rely on CAM-based methods. The heatmaps generated by these methods only reflect the contribution of certain image regions to classification tasks or the relevance of text features, failing to capture the role of multimodal features in recommender systems. Most existing multimodal recommendation works use use t-SNE (tsne, ) to analyze the modality features because t-SNE can visualize multimodal feature distributions. But t-SNE can only show the distribution of features in 2-dimensional space, which lacks the intuitive effectiveness of CAM-generated heatmaps and is not suitable for understanding and analyzing specific content features. Consequently, existing attribution analysis methods exhibit significant deficiencies in analyzing multimodal features within multimodal recommender systems.
2.3. Parameter-Efficient Adaptation
In the field of computer vision (CV) and natural language processing (NLP), several studies have identified that semantic differences can restrict the expressiveness of features, thereby impacting the model’s performance as well as downstream tasks (nlp1, ; cv1, ). To address this challenge, researchers have explored techniques such as fine-tuning and adapters to bridge these semantic differences (nlp2, ; softprompt, ; lora, ). However, in the context of recommendation scenarios where data is continuously changing and expanding, performing full fine-tuning of the encoder poses significant challenges. It often requires a significant amount of time and computational resources, and it may even lead to a decline in the model’s ability to generalize to new data. Given these practical limitations, retraining the feature encoder from scratch for every change in the data is unrealistic. Most of the existing multimodal recommendation methods directly use pre-trained features without fine-tuning, but the content feature extracted by the pre-trained feature encoders may not be sufficient to reflect the users’ preferences, so we consider introducing feature adapters. One such approach is Low-Rank Adaptation(LoRA) (lora, ), which reduces the number of trainable parameters by integrating trainable low-rank decomposition matrices into the Transformer architecture. Another approach is Prompt Tuning (softprompt, ), which offers a different perspective by incorporating learnable embedding vectors, acting as hints, into the inputs of the pre-trained model.
However, adding adapters to the middle layer of the pre-trained model may lead to some information loss. Existing parameter tuning methods are not generalized for recommendation tasks, as they fail to adequately incorporate behavioral information, which intuitively reflects the users’ preferences and is crucial for an effective recommender systems.
3. PRELIMINARIES
In this section, we analyze and prove in detail the problems of the content features extracted by pre-trained encoders for recommendation tasks. In section 3.1 we introduce an attribution analysis for visual attribution analysis of content features in multimodal recommender systems. In section 3.2, we specifically analyze the deficiencies of pre-trained feature encoders. In section 3.3 we analyze and prove the problems from a mathematical and theoretical point of view.
3.1. Attribution Analysis
Existing methods for analyzing the weights of image features are predominantly based on class activation mapping (CAM) (CAM, ). However, the heatmaps generated by CAM only reflect the contribution of certain image regions to the classification task and fail to assess their impact on recommendation models (scoreCAM, ; gScoreCAM, ). These limitations restrict our in-depth understanding of modality feature processing in recommender systems. Thus we designed the attribution analysis method for analyzing image features in multimodal recommendations based on the characteristics of recommender systems.
To visually illustrate the issues raised, we introduce a intuitively analysis for the attribution of multimodal features in the recommender system. This method aims to explore the effectiveness of content features extracted by pre-trained multimodal feature encoders in multimodal recommender systems. It also provides a visual demonstration of the effectiveness of our proposed adapter in bridging the gap between the content features extracted by the pre-trained encoders and the features required by the recommender system. To accurately identify which parts of the content features users pay attention to, we use cosine similarity as a measurement. By calculating the cosine similarity of each pixel to the behavioral features in the recommendation models, we generate a saliency heatmap. This heatmap is created by weighting and summing the masked feature maps based on the calculated similarities. This approach visualizes the portions of the content features that are truly effective in the recommendation task, providing a deeper understanding of the role of multimodal features in recommender systems.
(1) $\text{similarity}(x,y)=\frac{{v_{x}\cdot v_{y}}}{{\|v_{x}\|\cdot\|v_{y}\|}}$
Similar to the concept behind CAM-based approaches, we choose the target layer for visualization to ensure interpretability (scoreCAM, ; gScoreCAM, ). In the case of image encoders, for both ResNet (resnet, ) and ViT (ViT, ) versions, we select the channels closest to the CLIP prediction layer. In convolutional neural networks (CNNs), as the layers go deeper, the information represented in the feature maps becomes more abstract and high-level. The feature maps of the last ReLU layer capture the most important high-level features in the input image. ReLU-activated feature maps contain only positive or zero activations, which helps to emphasize the regions that the model considers more relevant. For the ViT series, as the gradients are zero for all layers except for the [CLS] token and only the [CLS] token is used for final predictions in the last layer embedding, we choose the penultimate layer. Consequently for the ResNet series, we select the last ReLU layer of the final Bottleneck as the target layer, for the ViT series, we use the second-to-last ResidualAttentionBlock as the target layer.
In the first stage, we perform a forward pass of the product images through CLIP, obtaining $N$ channels at the target layer denoted as ${A_{i}}^{N}$. The purpose is to discriminate the contributions of different targets across different channels. Subsequently, for each channel, we upsample it, denoted as $A_{i}^{up}$. This is done to generate masked images corresponding to different parts and to calculate the similarity weights of different regions. In the second stage, element-wise multiplication is performed between each mask $A_{i}^{up}$ and all color channels $x$ of the input image, and then the result is compared with the corresponding to behavioral features to compute the similarity. We name our method Behavior Information CAM (BeCAM) and its saliency map is computed by:
(2) $M_{\mathrm{BeCAM}}^{i}=\sum_{i}^{N}A_{i}\times\text{similarity}((x\odot A_{i}^{up}),{ID}_{i})$
where $\odot$ is the Hadamard product.
3.2. Deficiency Analysis
Figure 4. An illustration of a theoretical analysis of deficiency analysis
Figure 5. Comparison of efficient parameter tuning methods. (a) Low-Rank Adaptation (b) Soft Prompt Turning (c) Our proposed adapter.
By analyzing the generated heatmap (shown in Figure 2), it is clear that the pre-trained feature encoder suffers from the issues of wrong region of interest and insufficient region of interest on the product dataset. This can be visualized by the fact that the content features extracted by the pre-trained feature encoders have the problems of information drift and information omission. This is a very visible demonstration of the low information value density of the multimedia content itself, resulting the content features are insufficient to reflect users’ preferences.
Specifically, as shown in Figure 2, although the user wants to buy clothes, the feature encoder incorrectly focuses on the model’s face, resulting in information drift. There are also products whose content features miss some key details of the product, such as only focusing on certain details of the clothes and failing to focus on its whole, resulting in information omission. Information drift can cause the recommender system to ignore useful and critical information. This leads to a loss of important guidance for the recommendation task. It also results in a mismatch between recommendations and users’ preferences. Information omission can result in important visual or textual cues being ignored or incomplete. This causes a loss of critical information needed for the recommendation task. These issues collectively compromise the overall effectiveness of the recommender system. As a result, directly applying the content features extracted by the pre-trained encoder in the multimodal recommender system will inevitably impair the performance of the recommender system.
3.3. Theoretical Analysis
In this section, we reveal the deficiencies of content features extracted by pre-trained feature encoders within the embedding space. We analyze the gap between these content features and the ideal features required by the recommendation model. Furthermore, we demonstrate that optimizing the consistency between the extracted features and the ideal features necessary for accurate recommendation predictions can significantly enhance recommendation performance. This analysis provides theoretical support for the effectiveness of our proposed adapter.
The embedding space captures the semantic relationships and similarities between features. Analyzing features in the embedding space enables a better understanding and interpretation of their relationships. Formally, considering a node ${i}$ and its representation $\boldsymbol{e_{i}}$ in the embedding space, which represents the ideal features required by the recommendation model. On the basis of the item feature representation $\boldsymbol{e_{i}}$ in the embedding space, the recommender system calculates the probability distribution of the rating or clicking behavior of the user $u$ on the item $i$, which can be viewed as the prediction $P(r_{ui}|\boldsymbol{e_{u}},\boldsymbol{e_{i}})$. Where $r_{ui}$ stands for the rating or click behavior of the user $u$ on item $i$, $\boldsymbol{e_{u}}$ represents the user’s feature, and $\boldsymbol{e_{i}}$ signifies the item’s feature. Our objective is to find the optimal user feature $e_{u}$ and item feature $\boldsymbol{e_{i}}$ to maximize the posterior distribution $P(r_{ui}|\boldsymbol{e_{u}},\boldsymbol{e_{i}})$. In other words, we aim to maximize the posterior distribution for the user feature $\boldsymbol{e_{u}}$ and item feature $\boldsymbol{e_{i}}$ parameters. This can be formally described by the following objective function:
(3) $L(\boldsymbol{e_{u}},\boldsymbol{e_{i}})=\arg\max_{\boldsymbol{x_{u}},\boldsymbol{e_{i}}}P(r_{ui}|\boldsymbol{e_{u}},\boldsymbol{e_{i}})$
The representation $\boldsymbol{e_{i}^{\prime}}$ extracted by the feature encoder can be seen as the prior distribution. However, due to semantic differences between the natural data and e-commerce data, there is inevitably some deviations between the features extracted by the pre-trained feature encoder and the ideal features. Let us denote this deviation as $\theta$.
(4) $\theta=\arccos\left(\frac{\boldsymbol{e_{i}}\cdot\boldsymbol{e_{i}^{\prime}}}{\|\boldsymbol{e_{i}}\|\|\boldsymbol{e_{i}^{\prime}}\|}\right)$
Analyzing $\theta$ helps understand the deviation between the features extracted by the feature encoder and the ideal features. A smaller $\theta$ indicates that the two representations are very close in the embedding space. This implies a high consistency between the features extracted by the feature encoder and the ideal features, which benefits the recommendation system in accurately capturing user interests and item similarities. By reducing $\theta$, the features become closer to the ideal features, thereby enhancing the performance of the recommendation system.
The part $p$ of the representation $\boldsymbol{e_{i}^{\prime}}$ extracted by the encoder that is truly effective for the recommendation task can be represented as follows:
(5) $p=\left(\boldsymbol{e_{i}^{\prime}}\cdot\frac{\boldsymbol{e_{i}}}{\|\boldsymbol{e_{i}}\|}\right)\frac{\boldsymbol{e_{i}}}{\|\boldsymbol{e_{i}}\|}=\boldsymbol{e_{i}^{\prime}}\cdot\cos(\theta)$
When the deviation $\theta$ is large, as $\boldsymbol{e_{i}^{\prime}}$ in Figure 4, content features exhibit the issues of information omission, causing the effective length of $p$ deviates significantly from $\boldsymbol{e_{i}}$. This results in a lack of crucial information in the extracted features, reflected as insufficient region of interest in the heatmap. Conversely, when $\theta$ is too large,as $\boldsymbol{e_{j}}$ in Figure 4, content features exhibit the issues of information drift, with $p$ being located in the wrong quadrant and its effective direction opposing $\boldsymbol{e_{i}}$. This introduces a large amount of incorrect information, reflected as incorrect region of interest in the heatmap.
In such circumstances, the recommendation system is easily to make incorrect recommendations, because the feature representation contradicts the users’ preferences, failing to appropriately reflect the characteristics of items or users. In the recommendation task, we aim for a smaller deviation between $\boldsymbol{e_{i}^{\prime}}$ and $\boldsymbol{e_{i}}$. To formalize this objective, we define an error function $f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})$ to measure the gap between the representations:
(6) $f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})=\left(1-\frac{{\boldsymbol{e_{i}^{\prime}}\cdot\boldsymbol{e_{i}}}}{{\|\boldsymbol{e_{i}^{\prime}}\|\|\boldsymbol{e_{i}}\|}}\right)$
We seek to minimize the expected deviation measure $\Delta$. We can define $\Delta$ as:
(7) $\Delta=\mathbb{E}_{P(\boldsymbol{e_{i}^{\prime}})}[f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})]$
The function $f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})$ measures the deviation between representations. In recommendation tasks, the quality of the feature representation $\boldsymbol{e_{i}^{\prime}}$ directly affects the accuracy and effectiveness of recommendations. A decline in the quality of $\boldsymbol{e_{i}^{\prime}}$, indicated by an increase in the value of the error function $f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})$, donates an augmented diversity between the representation extracted by the pre-trained encoder $\boldsymbol{e_{i}^{\prime}}$ and the ideal representation $\boldsymbol{e_{i}}$. Consequently, this leads to an increase in the expected deviation $\Delta$, subsequently influencing the posterior distribution $P(r_{ui}|\boldsymbol{e_{u},e_{i}})$. Such circumstances may impair the recommendation system’s ability to accurately capturing the users’ preferences, thereby reducing the accuracy and efficacy of recommendations. By minimizing $\Delta$, we can ensure that the representation $\boldsymbol{e_{i}^{\prime}}$ are closer to the ideal representation $\boldsymbol{e_{i}}$ in expectation. This adaptation of content features enhances the consistency with the ideal representation and thus improves the performance of the recommender system.
Figure 6. Visual analysis of the effect of adapter on feature purification. (a) is the original image. (b) is the similarity heatmap of the visual features extracted by the pre-trained CLIP. (c) (d) (e) are the similarity heatmaps of the visual features adapted by BeFA applied on FREEDOM, BM3 and LATTICE respectively
Table 1. Statistics of the experimental datasets
Dataset #User #Item #Behavior Density
TMALL 13,104 7,848 151,928 0.148%
Microlens 46,420 14,079 332,730 0.051%
H&M 43,543 16,915 369,945 0.050%
4. feature adapter
To address the discovered issues, we proposed Behavior-driven Feature Adapter (BeFA). This is a adapter for adapting multimodal content features that are employed by multimodal recommender systems. Specifically, the content features are first extracted by a pre-trained encoder, and then the content features are adapted by BeFA, and then fed into the multimodal recommender system for item and user modeling. BeFA and the downstream recommender system share the optimization objective and use End-to-End training.
4.1. Problem Formulation
Let $u\in\mathcal{U}$ and $i\in\mathcal{I}$ denote the user and item, respectively. The input behavioral features for user $u$ and item $i$ are represented as $\mathrm{E_{id}}\in\mathbb{R}^{d\times(|U|+|I|)}$, where $d$ is the embedding dimension. Each item modality feature is denoted as $\mathrm{E_{i,m}}\in\mathbb{R}^{d_{m}\times|I|}$, where $d_{m}$ is the dimension of the features, $m\in\mathcal{M}$ represents the modality, and $\mathcal{M}$ is the set of modalities. In this paper, we primarily focus on visual and textual modalities, denoted by =$\mathcal{M}=\{v,t\}$. It’s worth noting that our approach is adaptable to incorporate more than two modalities.
Next, the users’ historical behavior feature is represented as $R\in\{0,1\}^{|U|\times|I|}$, where each entry $R_{u,i}=1$ if user $u$ clicked item $i$, otherwise $R_{u,i}=0$ This historical interaction data $R$ can naturally be interpreted as a user-item bipartite graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$. In $\mathcal{G}$ , the vertex set $\mathcal{V}=\mathcal{U}\cup\mathcal{I}$ donates all items and users. The edge set =$\mathcal{E}=\{(u,i)|u\in\mathcal{U},i\in\mathcal{I},R_{u,i}=1\}$ represents the interacted users and items. The objective of multimedia recommendation is to accurately predict users’ preferences by ranking items for each user based on predicted preference scores $\hat{y}_{u,i}$.
Table 2. Performance Comparison on Different Recommender Models. The t-tests validate the significance of performance improvements with p-value $\leq$ 0.05.
TMALL Microlens H&M
Encoder Datasets R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20
BM3 0.0189 0.0298 0.0102 0.0132 0.0510 0.0851 0.0278 0.0375 0.0204 0.0320 0.0114 0.0144
BM3+BeFA 0.0212 0.0319 0.0115 0.0144 0.0566 0.0911 0.0314 0.0412 0.0266 0.0391 0.0157 0.0190
Imporve 12.17% 7.05% 12.75% 9.09% 10.98% 7.05% 12.95% 9.87% 30.39% 22.19% 37.72% 31.94%
LATTICE 0.0238 0.0356 0.0134 0.0167 0.0553 0.0886 0.0308 0.0402 0.0289 0.0427 0.0161 0.0197
LATTICE+BeFA 0.0260 0.0403 0.0183 0.0183 0.0593 0.0943 0.0328 0.0427 0.0317 0.0498 0.0171 0.0217
Imporve 9.24% 13.20% 36.57% 9.58% 7.23% 6.43% 6.49% 6.22% 9.69% 16.63% 6.21% 10.15%
FREEDOM 0.0212 0.0340 0.0113 0.0148 0.0474 0.0774 0.0262 0.0348 0.0348 0.0526 0.0188 0.0234
FREEDOM+BeFA 0.0253 0.0375 0.0136 0.0170 0.0503 0.0814 0.0279 0.0368 0.0409 0.0583 0.0226 0.0271
Improve 19.34% 10.29% 20.35% 14.86% 6.12% 5.17% 6.49% 5.75% 17.53% 10.84% 20.21% 15.81%
MGCN 0.0249 0.0380 0.0135 0.0171 0.0618 0.0972 0.0342 0.0442 0.0367 0.0549 0.0204 0.0251
MGCN+BeFA 0.0261 0.0395 0.0142 0.0179 0.0630 0.1000 0.0351 0.0456 0.0405 0.0594 0.0225 0.0274
Improve 4.82% 3.95% 5.19% 4.68% 1.94% 2.88% 2.63% 3.17% 10.35% 8.20% 10.29% 9.16%
CLIP Avg Imporve 11.39% 8.62% 18.71% 9.55% 6.57% 5.38% 7.14% 6.25% 16.99% 14.46% 18.61% 16.77%
BM3 0.0184 0.0299 0.0097 0.0129 0.0508 0.0842 0.0279 0.0373 0.0195 0.0304 0.0107 0.0135
BM3+BeFA 0.0224 0.0322 0.0125 0.0152 0.0537 0.0877 0.0299 0.0395 0.0248 0.0378 0.0149 0.0183
Imporve 21.74% 7.69% 28.87% 17.83% 5.71% 4.16% 7.17% 5.90% 27.18% 24.34% 39.25% 35.56%
LATTICE 0.0252 0.0374 0.0139 0.0173 0.0580 0.0953 0.0320 0.0426 0.0293 0.0439 0.0164 0.0202
LATTICE+BeFA 0.0266 0.0403 0.0147 0.0184 0.0633 0.1021 0.0340 0.0451 0.0316 0.0498 0.0171 0.0218
Imporve 5.56% 7.75% 5.76% 6.36% 9.14% 7.14% 6.25% 5.87% 7.85% 13.44% 4.27% 7.92%
FREEDOM 0.0197 0.0319 0.0107 0.0140 0.0613 0.0976 0.0337 0.0440 0.0364 0.0553 0.0192 0.0241
FREEDOM+BeFA 0.0259 0.0377 0.0141 0.0174 0.0641 0.1004 0.0356 0.0459 0.0417 0.0613 0.0234 0.0285
Improve 31.47% 18.18% 31.78% 24.29% 4.57% 2.87% 5.64% 4.32% 14.56% 10.85% 21.88% 18.26%
MGCN 0.0266 0.0403 0.0144 0.0182 0.0693 0.1075 0.0388 0.0496 0.0405 0.0613 0.0218 0.0272
MGCN+BeFA 0.0275 0.0414 0.0152 0.0185 0.0702 0.1085 0.0389 0.0498 0.0451 0.0655 0.0246 0.0299
Improve 3.38% 2.73% 5.56% 1.65% 1.30% 0.93% 0.26% 0.40% 11.36% 6.85% 12.84% 9.93%
ImageBind Avg Imporve 15.54% 9.09% 17.99% 12.53% 5.18% 3.77% 4.83% 4.12% 15.24% 13.87% 19.56% 17.92%
4.2. Behavior-driven Feature Adapter
Due to the shortcomings of pre-trained feature encoders, their extracted content features contain a significant amount of irrelevant and erroneous information. To better utilize modality information,we propose Behaviour-driven Feature Adapter(BeFA) for adapting content features. Firstly, we decouple the original item content features $\mathrm{E_{i,m}}$ into features in the decoupled feature space $\mathrm{\dot{E}_{i,m}}$:
(8) $\dot{\mathrm{E}}_{i,m}=\mathrm{W}_{1}\mathrm{E}_{i,m}+\mathrm{b}_{1},$
where $\mathbf{W}_{1}\in\mathbb{R}^{d\times d_{\lambda}}$ and $\mathbf{b}_{1}\in\mathbb{R}^{d_{m}}$ represent trainable transformation matrix and bias vector, respectively.
Considering that the behavioral information fully reflects the users’ preference, we filter the preference-related content features with the help of behavioral information guidance:
(9) $\ddot{\mathrm{E}}_{i,m}=f_{gate}^{m}(\mathrm{E}_{i,id},\dot{\mathrm{E}}_{i,m})=\mathrm{E}_{i,id}\odot\sigma(\mathrm{W}_{2}\dot{\mathrm{E}}_{i,m}+\mathrm{b}_{2}),$
where $\mathbf{W}_{2}\in\mathbb{R}^{d_{\lambda}\times d_{\lambda}}$ and $\mathbf{b}_{2}\in\mathbb{R}^{d_{m}}$ are trainable parameters. Here, $\odot$ denotes element-wise multiplication, and $\sigma$ is the Tanh non-linear transformation.
Finally, the decoupled content features are recombined. Given that behavioral information reflects user preferences, we determine the weights of different decoupled content features based on the guidance provided by this behavioral information:
(10) $\overline{\mathrm{E}}_{i,m}=f_{merge}^{m}(\mathrm{E}_{i,id},\ddot{\mathrm{E}}_{i,m})=\mathrm{E}_{i,id}\odot\sigma(\mathrm{W}_{3}\ddot{\mathrm{E}}_{i,m}+\mathrm{b}_{3}),$
where $\mathbf{W}_{3}\in\mathbb{R}^{d_{\lambda}\times d}$ and $\mathbf{b}_{3}\in\mathbb{R}^{d}$ are trainable parameters. Similarly, $\odot$ denotes element-wise multiplication, and $\sigma$ is the sigmoid non-linear transformation. It is important to note that although Equations 9 and 10 are formally similar, they achieve different effects.
Moreover, we introduce ReLU activation functions and Dropout between each transformation layer. This enables the model to learn non-linear relationships, thereby improving its fitting ability and expressive power. Additionally, it enhances the model’s generalization, making it more suitable for practical application scenarios.
5. EXPERIMENTS
In this section, we conduct extensive experiments to evaluate the performance of the proposed adapter on three public datasets to answer following research questions.
  • RQ1: How effective is BeFA applied to existing multimodal recommender systems?
  • RQ2: How effective is BeFA in adapting content features?
  • RQ3: Why purifying modality information can achieve better recommendation performance?
  • RQ4: How do hyper-parameter settings impact the performance of BeFA?
5.1. Experimental Settings
5.1.1. Datasets
We conducted experiments on three publicly available datasets: (a) TMALL [^1] https://tianchi.aliyun.com/dataset/140281; (b) Microlens [^2] https://recsys.westlake.edu.cn/MicroLens-50k-Dataset/; and (c) H&M [^3] https://www.kaggle.com/datasets/odins0n/handm-dataset-128x128. We performed 10-core filtering on the raw data. The detailed information of the filtered data is presented in Table 1. For multimodal information, we utilized pre-trained CLIP (clip, ) and ImageBind (imagebind, ) to generate aligned multimodal features.
5.1.2. Comparative Model Evaluation
To verify the prevalence of feature encoder defects, we used ViT-B/32-based CLIP (clip, ) and ImageBind (imagebind, ) for our experiments. CLIP is a general-purpose feature encoder capable of handling a wide range of data types and application domains. It realizes cross-modality feature representations and exhibits highly efficient performance, making it widely adopted in practical applications and performing well across various tasks. ImageBind is the latest cross-modality feature encoder, which can handle a wider variety of modality information, and achieves a new SOTA performance in the emerging zero-shot recognition task for various modalities, even outperforming the previous specialized models trained specifically for these modalities.
To evaluate the effectiveness of our proposed adapter, we applied it to several representative multimodal recommendation models, including LATTICE (lattice, ), BM3 (bm3, ), FREEDOM (FREEDOM, ), and MGCN (mgcn, ) for comparison. Additionally, we compared our adapter with existing efficient parameter tuning methods such as LoRA (lora, ) and Soft-Prompt Tuning (softprompt, ).
Table 3. Performance Comparison with other Efficient Parameter Adaptation methods. The best result is in boldface and the second best is underlined. The t-tests validate the significance of performance improvements with p-value $\leq$ 0.05.
TMALL Microlens H&M
Dadasets R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20
[…]
BM3+LoRA 0.0215 0.0333 0.0117 0.0150 0.0510 0.0850 0.0280 0.0376 0.0191 0.0290 0.0110 0.0135
[…]
FREEDOM 0.0212 0.0340 0.0113 0.0148 0.0474 0.0774 0.0262 0.0348 0.0348 0.0526 0.0188 0.0234
[…]
FREEDOM+BeFA 0.0243 0.0364 0.0131 0.0164 0.0503 0.0814 0.0279 0.0368 0.0409 0.0583 0.0226 0.0271
[…]
MGCN+LoRA 0.0260 0.0391 0.0141 0.0179 0.0598 0.0963 0.0335 0.0438 0.0367 0.0554 0.0203 0.0252
[…]
Avg Improve 2.44% 1.71% 7.89% 0.48% 6.08% 4.98% 6.25% 5.67% 11.11% 9.59% 12.58% 11.44%
i) Multimodal Recommendation Models:
  • LATTICE (lattice, ): This method designs a new modality-aware structure learning layer that learns the item-item structure of each modality and aggregates multiple modalities to obtain latent item graphs. Based on the learned latent graphs, graph convolution is performed to explicitly inject higher-order item affinities into item representations.
  • BM3 (bm3, ): This method proposes a simple self-supervised learning framework that generates target views of users or items for comparison learning without negative samples.
  • FREEDOM (FREEDOM, ): This method devises a degree-sensitive edge pruning method to denoise user-item interaction graphs, rejecting potentially noisy edges with high probability when sampling the graph.
  • MGCN (mgcn, ): This method introduces a behavior-aware fuser to effectively model user preferences by dynamically learning the relative significance of various modality features. Additionally, it integrates the fuser with a self-supervised auxiliary task aimed at maximizing the mutual information between the fused multimodal features and behavioral features. This approach enables the concurrent capture of complementary and supplementary preference information.
ii) Efficient Parameter Tuning Methods
  • Low-Rank Adaptation (LoRA) (lora, ): This method aims to enhance the representation of a model while keeping the number of model parameters relatively small. It provides additional learning power by introducing lower rank matrices to linearly transform the hidden representation of the original model. This approach strikes a balance between parameter efficiency and model performance.
  • Soft Prompt Turning (softprompt, ): This method enhances model performance by fine-tuning small, task-specific embeddings (prompts) inserted into the input, thereby adapting pre-trained language models to new tasks without altering their original weights.
5.1.3. Evaluation Protocols
For a fair comparison, we follow the evaluation settings in (lattice, ; bm3, ) with the same 8:1:1 data splitting strategy for training, validation and testing. Besides, we follow the all-ranking protocol to evaluate the top-K recommendation performance and report the average metrics for all users in the test set: $\mathrm{R@K}$ and $\mathrm{N@K}$, which are abbreviations for $\mathrm{Recall@K}$ (recall, ) and $\mathrm{NDCG@k}$ (NDCG, ), respectively.
5.1.4. Implementation Details
We implement MMRec [^4] https://github.com/enoche/MMRec (mmrec, ) based on PyTorch, which is a unified public repository designed for multimodal recommendation methods. To ensure fair comparison, we employed the Adam optimizer to optimize all methods and referred to the best hyperparameter settings reported in the original baseline paper. For general settings, we initialized embeddings with Xavier initialization of dimension 64, set the regularization coefficient to $\lambda_{E}=10^{-4}$, and batch size to $B=2048$. Early stopping and total epochs are fixed at 10 and 1000, respectively. We selecte the best model with the highest $\mathrm{Recall@20}$ metric on the validation set and reported metrics on the test set accordingly.
5.2. Overall performance(RQ1)
5.2.1. Effectiveness
(a) Comparison of model performance before and after applying BeFA
Table 2 shows the performance comparison of BeFA applied to existing multimodal recommendation models on three datasets. From the table,we find several observations:
  • Our adapter significantly improved recommendation performance on all three datasets. Specifically, our adapter has achieved an average improvement of 9.07% and 11.02% over the baseline in terms of $\mathrm{Recall@20}$ and $\mathrm{NDCG@20}$. This suggests that the content features extracted by the pre-trained encoder do have deficiencies that affect recommendation performance. Our processing of content features is very efficient and effectively reduces modality noise. This helps the multimodal recommendation models to better utilize the modality information and thus improve the recommendation performance.
  • Our adapter has significant performance improvements for all encoders. All pre-trained feature encoders are generally defective. We use CLIP and ImageBind as an example, and applying our adapter to the multimodal features extracted by these pre-trained feature encoders has led to a significant improvement in the performance of the recommender system. This suggests that the deficiencies are mainly due to the pre-training data, indicating that the content features extracted by the pre-trained encoders generally do not reflect the users’ preferences.
  • More powerful encoder can extract more sufficient content information. ImageBind has better overall preformance than CLIP when applied to recommendation models. This indicates that more advanced encoders are able to extract more sufficient content information. Thus better modelling of users and items to improve the performance of recommendations.
(b) Performance comparison with existing parameter tuning methods
Table 4. The training cost. #Param: number of tunable parameters, Time/E: averaged training time for one epoch, ’s’ means seconds. The results are collected on a GeForce RTX 4090 GPU.
Dataset Method #Param. Time/E
TMALL BM3 9.45M 0.38s
BM3+LoRA +4.10K +0.04s
BM3+SoftPrompt +0.13K +0.02s
BM3+BeFA +0.20M +0.08s
Microlens BM3 18.36M 0.98s
BM3+LoRA +4.10K +0.07s
BM3+SoftPrompt +0.13K +0.04s
BM3+BeFA +0.20M +0.15s
H&M BM3 21.26M 1.20s
BM3+LoRA +4.10K +0.04s
BM3+SoftPrompt +0.13K +0.01s
BM3+BeFA +0.20M +0.12s
Figure 7. The effect of LoRA and SoftPrompt Turning on the adaptation of features. (a) is the original heatmap, (b) is the heatmap of visual features adapted by LoRA, (c) is the heatmap of visual features adapted by SoftPrompt Turning and (d) is the heatmap of visual features adapted by BeFA.
Table 3 shows the comparison between our adapter and existing efficient parameter adaptation methods. From the table,we find several observations:
  • The performance of our adapter is overall significantly better than existing efficient parameter tuning methods. This indicates that our adapter is better suited for the recommendation task than generalized efficient parameter tuning methods. It suggests that our adapter employs behavioral information guidance to more effectively obtain content features related to users’ preferences, thereby improving the effectiveness of recommendations.
  • Existing parameter tuning approaches are not generalized for recommendation tasks. The performance decreases in some cases instead, which indicates that these existing generalized parameter tuning methods are not necessarily suitable for the recommendation tasks. This highlights the importance of incorporating behavioral information guidance in recommendation tasks to improve model performance.
Figure 8. The gaps in the effectiveness of our adapter on different datasets. (a) is the original image. (b) is the heatmap of the visual features extracted by the pre-trained CLIP. (c) is the heatmap of visual features adapted by BeFA.
[…]
HM_Microlens
Figure 9. The Distribution of representations in different multimodal recommendation model in visual modalities before and after adaptation by BeFA.
5.2.2. Efficiency
We also analyze the efficiency of our adapter with other existing efficient parameter tuning methods on the BM3 model. The specific model parameter counts and training costs are shown in Table 4. Compared to the overall model training time, our training time increases by only about 15%, indicating that our method is not only effective but also efficient.
Although the design of our adapter is more complex compared to LoRA and Soft-Prompt Turning the increase in the number of parameters after using our adapter is still small compared to the overall parameters of the recommender system itself, ranging from 0.93% to 2.09%. In terms of training costs, the increase in overall training time after using our adapter is still within an acceptable range. The training time for the recommender system using our adapter is comparable to that of using LoRA and Soft-Prompt Tuning. This suggests that although our adapter introduces additional parameters, it does not impose a significant burden on the training process. The substantial improvement in performance of our adapter highlights its efficiency, trading a small number of parameters and training time gains for a significant improvement in recommendation accuracy.
5.3. Visualization Analysis(RQ2&RQ3)
5.3.1. Case Study
(a)Effectiveness with different models
To visually demonstrate the effect of our adapter on content features, we applied it to the proposed visualization attribution analysis. Specifically, before calculating the similarity between content features and behavioral features in the attribution analysis, we processed the content features with our adapter. We then computed similarity weights with the behavioral features and generated new heatmaps using these newly calculated weights with the corresponding masked images.
As shown in Figure 6, by analysing the heatmap before and after our adapter, we can clearly find that our adapter’s enhancement of content features is very obvious. Our adapter significantly improves the attention areas of features extracted by the pre-trained feature encoder, addressing issues of information drift and information omission. The adapted features focus more on the recommended products, with a marked decrease in attention to background elements unrelated to the recommendation. This reduction in noise enhances feature representation and consequently improves recommendation performance.
Additionally, the adapted features more comprehensively capture product details, significantly increasing the coverage of the area of interest for products. This comprehensive extraction of product details not only improves feature recognition but also enhances the recommender system’s ability to model product features. This means that the recommender system is able to more accurately capture the core features of the product, thus providing more relevant and accurate results when making recommendations. Overall,our adapter enables multimodal recommender systems to better utilize content features by using behavioral information to guide feature adaptation, thus significantly improving recommendation performance.
(b) Effectiveness with different encoders
We utilize the content features extracted by two representative multimodal feature encoders, the ViT-32/B version of CLIP and ImageBind, and analyze the heatmaps of the content features within different downstream recommendation models after applying BeFA. Our analysis reveals that the content features extracted by both encoders exhibit issues of information drift and information omission. After applying BeFA, these content features are better focused on the recommended products themselves. This demonstrates that the deficiencies of pre-trained feature encoders are common in multimodal recommendation tasks, while our adapter effectively adapts to the content features extracted by different feature encoders, highlighting its general applicability.
(c) Comparison with existing adaptation methods
We visualize and analyze the content features adapted by LoRA and Soft-Prompt Tuning, as shown in Figure 7. The results indicate that while LoRA and Soft-Prompt Tuning provide some optimization, the degree of adaptation is limited. The adapted content features still exhibit issues such as content error and content insufficiency, failing to focus accurately on the content of the product itself. These methods fail to adequately address the issues of information drift and information omission. This is the reason for the overall poor performance of existing adaptation methods in recommendation tasks. In contrast, the impact of BeFA on content features content features is significantly better than that of existing methods, which is why BeFA performs well in the recommendation task.
(d) Analysis in different scenarios
We employed three datasets ranging from simple to complex scenarios. The H&M dataset comprises only product images with minimal interference, making it relatively clean. In contrast, the Microlens dataset comprises more complex images with more background interference. Our adapter demonstrated significant improvements across both datasets. Specifically, on the H&M dataset, it increases $\mathrm{Recall@20}$ and $\mathrm{NDCG@20}$ by 14.17% and 17.34%, respectively. Meanwhile, on the Microlens dataset, our adapter increases $\mathrm{Recall@20}$ and $\mathrm{NDCG@20}$ by 4.58% and 5.19%, respectively. These findings indicate the adapter’s effectiveness across varying levels of scene complexity. Moreover, as illustrated in Figure 8, even in complex scenarios, our adapter effectively removes irrelevant information and locates the recommended products as precisely as possible
5.3.2. Representation distribution analysis
To visually demonstrate the impact of BeFA on adjusting content features, we visualize the distribution of visual representations in the TMALL dataset. Figure 9 illustrates the comparison of our adapter’s application to feature distributions across various recommendation models. Specifically, we randomly sample 500 item representations from the TMALL dataset and map them to two-dimensional normalized vectors on the unit hypersphere $\mathcal{S}^{1}$ (i.e., a circle with a radius of 1) by using t-SNE (tsne, ). Subsequently, we employ Gaussian kernel density estimation(KDE) (KDE, ) to plot the distribution of these two-dimensional features.
Further examining the changes in the distribution of two dimensional features across different recommendation models before and after applying BeFA, we observe that the original feature distribution is more dispersed, forming multiple small dense regions. This distribution indicates that there is a large amount of noise in the visual feature extraction process, hindering the models’ ability to effectively distinguish key features of different products. In contrast, after applying BeFA, the feature distribution becomes more uniform and smooth, which indicates that our adapter has significant effects in eliminating noise. This improvement not only enhances feature recognition capability but also improves the model’s ability to distinguish different content features (pre, ).
Moreover, comparing the changes in feature distributions across different models, it can be found that BeFA has strong adaptability. The feature distributions of various multimodal recommendation models show significant improvements after applying BeFA. This indicates that our adapter is not only effective on a single model but also has strong generalization capabilities across different recommender systems. This adaptability highlights BeFA’s wide potential and value in practical applications.
5.4. Sensitivity Analysis(RQ4)
Figure 10. Performance comparison w.r.t. different size of the decoupling space relative to the original embedding space $\lambda$.
[…]
We investigate the size of the decoupling space relative to the original embedding space, as shown in Fig.10. The results show that the optimal size of the decoupling space is about four times the size of the original embedding space on some datasets. When the decoupling space dimension is small (e.g., 0.125x and 0.25x), it may create an information bottleneck, leading to poor performance across all datasets in both $\mathrm{Recall@20}$ and $\mathrm{NCGD@20}$. A smaller decoupling space cannot hold sufficient information, making it difficult for the model to accurately capture and differentiate features, thereby negatively impacting recommendation performance. Conversely, an excessively large decoupling space dimension may lead to dimensionality catastrophe, introducing noise and redundant features which complicates the model’s ability to effectively distinguish and learn features in a high-dimensional space. Additionally, its number of parameters and the computational cost will increase dramatically, bringing additional burden to the training process, which is impractical for practical application scenarios.
Furthermore, we find that there is a significant difference in the performance of Microlens, TMALL and H&M datasets under varying decoupling space sizes. For the Microlens dataset, both $\mathrm{Recall@20}$ and $\mathrm{NCGD@20}$ achieve the optimal performance when the decoupling space size is four times of the original embedding space. In contrast, in the TMALL dataset, although the overall trend is similar, the optimal point fluctuates slightly, suggesting that the size of the decoupling space may need to be adjusted on different datasets to obtain optimal results. The overall adapter performance fluctuates greatly with the change in the decoupling space size, highlighting the challenge of finding the optimal size. This is one of the limitations of our work, which could be addressed by considering the decoupling space size as a learnable parameter in future research. This approach could adaptively adjust the decoupling space size for different application scenarios to achieve the best performance.
6. Conclusion
In this paper, we introduce an attribution analysis for visually analyzing deficiencies of the content features. We found that the content features of certain products suffer from information drift and information omission, and theoretically prove that these deficiencies lead to a decrease in the performance of the recommender systems. To address these issues, we propose Behavior-driven Feature Adapter (BeFA) adapter which adapts content features through the guidance of behavioral information, thereby enhancing the expressiveness of the content features and significantly improve the performance of the recommender systems
In our future work, we consider introducing the size of the decoupling space as a learnable parameter into the training of the adapter, so that the optimized performance can be achieved in different scenarios, enhancing its general applicability.

### v2

@@
 BeFA: A General Behavior-driven Feature Adapter for Multimedia Recommendation
-Abstract.
+Abstract
 Multimedia recommender systems focus on utilizing behavioral information and content information to model user preferences.
@@
 To verify our hypothesis, we introduce an attribution analysis method for visually and intuitively analyzing the content features.
-The results indicate that certain products’ content features exhibit the issues of information drift and information omission, reducing the expressive ability of features.
+The results indicate that certain items’ content features exhibit the issues of information drift and information omission, reducing the expressive ability of features.
 Building upon this finding, we propose an effective and efficient general Behavior-driven Feature Adapter (BeFA) to tackle these issues.
@@
 Extensive experiments demonstrate the effectiveness of the adapter across all multimedia recommendation methods.
-The code will be publicly available upon the paper’s acceptance.
-Multimodal Recommendation, Feature Adapter, Information Drift, Information Omission
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06submissionid: 123-A56-BU3
-Figure 1.
-Illustration of content features that do not accurately reflect the users’ preferences.
-Excessive irrelevant information hinders the recommender system’s ability to effectively model the users’ true preferences.
-1. Introduction
-Recommender systems have gained widespread adoption across various domains, aiming to assist users in discovering information that aligns with their preferences (vldb1, ; vldb2, ).
-In the case of multimedia platforms, the abundance of data resources provides recommender systems with increased opportunities to accurately model user preferences (survey, ).
-Existing multimedia recommendation methods typically involves two main steps (w2go, ; survey, ).
+Our code is made publicly available on https://github.com/fqldom/BeFA.
+Introduction
+Recommender systems have gained widespread adoption across various domains, aiming to assist users in discovering information that aligns with their preferences (Parchas et al.
+2020; Liu et al.
+2017).
+In the case of multimedia platforms, the abundance of data resources provides recommender systems with increased opportunities to accurately model user preferences (Zhou et al.
+2023a).
+Existing multimedia recommendation methods typically involves two main steps (Yuan et al.
+2023; Zhou et al.
+2023a).
 First, a pre-trained feature encoder is employed to capture content features from diverse modalities.
@@
 Recently, researchers have focused on enhancing the quality of these representations through self-supervised learning.
-For instance, SLMRec (slmrec, ) investigates potential relationships between modalities through the use of contrastive learning, thereby obtaining a powerful representation.
-BM3 (bm3, ) utilizes a dropout strategy to construct multiple views and reconstructs the interaction graph, incorporating intra- and inter-modality contrastive loss to facilitate effective representation learning.
-MICRO (micro, ) and MGCN (mgcn, ) maximize the mutual information between content features and behavioral features with a self-supervised auxiliary task, and have achieved excellent performance.
-Figure 2. Results of visualisation attribution analysis on the TMALL dataset.(a) is the original product image and (b) is its corresponding heatmap.
-Despite the good progress made by existing methods in utilizing content and behavioral information more effectively, a crucial yet easily overlooked problem arises: Are content features obtained from pre-trained encoders containing sufficient features to reflect user preferences?
-Intuitively, multimedia content inherently exhibits the characteristic of low informational value density, where a significant portion of the presented information may be irrelevant to the users’ focus (micro, ; survey, ).
+For instance, SLMRec (Tao et al.
+2023) investigates potential relationships between modalities through the use of contrastive learning, thereby obtaining a powerful representation.
+BM3 (Zhou et al.
+2023c) utilizes a dropout strategy to construct multiple views and reconstructs the interaction graph, incorporating intra- and inter-modality contrastive loss to facilitate effective representation learning.
+MICRO (Zhang et al.
+2022) and MGCN (Yu et al.
+2023) maximize the mutual information between content features and behavioral features with a self-supervised auxiliary task, and have achieved excellent performance.
+Figure 1: Illustration of content features that do not accurately reflect the users’ preferences.
+Excessive irrelevant information hinders the recommender system’s ability to effectively model the users’ true preferences.
+Figure 2: Results of visualisation attribution analysis on the TMALL dataset.
+The first row contains the original item images while the second row displays the corresponding heatmaps.
+The four samples on the left reflect information drift and the four samples on the right reflect information omission.
+Despite the good progress made in utilizing content and behavioral information more effectively, a crucial yet easily overlooked problem arises: Are content features obtained from pre-trained encoders containing sufficient features to reflect user preferences?
+Intuitively, multimedia content inherently exhibits the characteristic of low informational value density, where a significant portion of the presented information may be irrelevant to the users’ focus (Zhou et al.
+2023a).
 Pre-trained feature encoders extract information from the entire content simultaneously, which can result in content features that do not truly reflect the users’ preferences (as shown in Figure 1).
@@
 This method evaluates the extent to which content features can reflect user preferences, enabling researchers to visually assess the quality of content features for the first time.
-The results indicate that not all products’ content features accurately reflect user preferences.
-Due to the presence of irrelevant information, certain products’ content features exhibit the issues of information drift and information omission.
-As shown in Figure 2, some products’ content features do not include information about the products that users are interested in, but instead erroneously include information about unrelated items, a phenomenon we term information drift.
-There are also some products’ content features that omit certain key details of the products, which we refer to as information omission.
+The results indicate that not all items’ content features accurately reflect user preferences.
+Due to the presence of irrelevant information, certain items’ content features exhibit the issues of information drift and information omission.
+As shown in Figure 2, some items’ content features do not include information about the items that users are interested in, but instead erroneously include information about unrelated items, a phenomenon we term information drift.
+There are also some items’ content features that omit certain key details of the items, which we refer to as information omission.
 These issues ultimately prevent recommender systems from accurately modeling user preferences.
-To verify the rationality of these findings, we provide a theoretical analysis and explanation.
 Furthermore, we propose a plug-and-play general Behavior-driven Feature Adapter (BeFA) to address the discovered issues.
 This adapter effectively decouples, filters and reconstructs content features, leveraging behavioral information as a guide to obtain more precise representations of content information.
-Extensive experiments demonstrate the adapter’s effectiveness across various multimedia recommendation methods and feature encoders.
+Extensive experiments demonstrate the adapter’s effectiveness across various recommendation methods and feature encoders.
 Our main contributions can be summarized as follows:
   • We introduce a similarity-based visual attribution method, which enables researchers to visually analyze the quality of content features for the first time.
-  • We experimentally revealed the issues of information drift and information omission in content features. We also provide a theoretical analysis to validate the rationality of these findings.
+  • We experimentally revealed the issues of information drift and information omission in content features.
   • We propose a general behavior-driven feature adapter, which obtain more precise content representation through decoupling and reconstructing content features.
-Figure 3. Pipeline of the proposed attribution analysis method.
-2. RELATED works
-2.1. Multimedia Recommendation
-Collaborative filtering (CF)-based approaches have achieved great success in recommender systems, relying on behavioral similarities to make top-k recommendations (cf1, ; cf2, ).
-Since user preferences are usually influenced by multimodal information, it facilitates researchers to integrate multimodal content information into CF-based approaches with the aim of improvement.
-Typically, content features are first extracted using pre-trained neural networks and then fused with behavioral features to enhance preference modeling (vbpr, ; w2go, ).
-Considering that behavioral information can be naturally modeled as a bipartite graph structure, researchers begin work on recommender systems based on Graph Convolutional Network(GCN) (gcn1, ; gcn2, ).
-For instance, MMGCN citemmgcn constructs modality-specific user-item interaction graphs to model user preferences specific to each modality.
-LATTICE (lattice, ) adds links between items with similar modality characteristics and create a separate item-item graph for each modality.
-FREEDOM (FREEDOM, ) constructs item-item graphs to help learn item semantic relations and reduces the graphs according to the sensitivity edge pruning technique.
-Meanwhile, researchers began to focus on enhancing the quality of representations through self-supervised learning.
-BM3 (bm3, ) learns user and item representations by reconstructing the user-item interaction graph and aligning modality features from both inter and intra-modality perspectives.
-MGCN (mgcn, ) comprehensively models user preferences by adaptively learning the relative importance of different modality features.
-However, existing methods usually directly utilize features extracted from pre-trained feature encoders as side information for each item.
-The pre-trained feature encoders are trained on large-scale datasets collected from the natural world.
-The features of natural data and e-commerce data are very different.
-In e-commerce scenarios, the content information often contains a lot of irrelevant information and exhibits the characteristic of low value density.
-For example, only a small portion of the content image contain relevant information, while the rest is considered noise.
-This disparity leads to a substantial amount of noise in the contetn features extracted by pre-trained encoders when applied to multimodal recommender systems.
-Consequently, the content features do not contain enough features to sufficiently reflect users’ preferences.
-Fusing these irrelevant content features with behavioral features may mislead user preference modeling, resulting in suboptimal recommendation performance.
-2.2. Attribution Analysis
-Methods for modality feature attribution analysis have been widely studied, aiding in the visualization of modality feature content.
-Class Activation Mapping (CAM) (CAM, ) is a method used to visualize the decision-making process of convolutional neural networks by generating heatmaps that highlight regions of an image contributing to a particular class prediction.
-Grad-CAM (gradcam, ) enhances this by using the gradient of the target class with respect to the feature maps of a specific convolutional layer, creating heatmaps that can be applied to various network architectures without altering the network structure.
-Score-CAM (scoreCAM, ) further improves upon Grad-CAM by weighting the activation map using output scores instead of gradient information, thus avoiding gradient-related problems.
-In multimodal recommender systems, t-SNE (tsne, ) is often used to visualize the distribution of samples in high-dimensional feature spaces, providing insights into the relationships and distributions among different modality features.
-However, current methods (cvpr, ; gScoreCAM, ) for analyzing image feature weights predominantly rely on CAM-based methods.
-The heatmaps generated by these methods only reflect the contribution of certain image regions to classification tasks or the relevance of text features, failing to capture the role of multimodal features in recommender systems.
-Most existing multimodal recommendation works use use t-SNE (tsne, ) to analyze the modality features because t-SNE can visualize multimodal feature distributions.
-But t-SNE can only show the distribution of features in 2-dimensional space, which lacks the intuitive effectiveness of CAM-generated heatmaps and is not suitable for understanding and analyzing specific content features.
-Consequently, existing attribution analysis methods exhibit significant deficiencies in analyzing multimodal features within multimodal recommender systems.
-2.3. Parameter-Efficient Adaptation
-In the field of computer vision (CV) and natural language processing (NLP), several studies have identified that semantic differences can restrict the expressiveness of features, thereby impacting the model’s performance as well as downstream tasks (nlp1, ; cv1, ).
-To address this challenge, researchers have explored techniques such as fine-tuning and adapters to bridge these semantic differences (nlp2, ; softprompt, ; lora, ).
-However, in the context of recommendation scenarios where data is continuously changing and expanding, performing full fine-tuning of the encoder poses significant challenges.
-It often requires a significant amount of time and computational resources, and it may even lead to a decline in the model’s ability to generalize to new data.
-Given these practical limitations, retraining the feature encoder from scratch for every change in the data is unrealistic.
-Most of the existing multimodal recommendation methods directly use pre-trained features without fine-tuning, but the content feature extracted by the pre-trained feature encoders may not be sufficient to reflect the users’ preferences, so we consider introducing feature adapters.
-One such approach is Low-Rank Adaptation(LoRA) (lora, ), which reduces the number of trainable parameters by integrating trainable low-rank decomposition matrices into the Transformer architecture.
-Another approach is Prompt Tuning (softprompt, ), which offers a different perspective by incorporating learnable embedding vectors, acting as hints, into the inputs of the pre-trained model.
-However, adding adapters to the middle layer of the pre-trained model may lead to some information loss.
-Existing parameter tuning methods are not generalized for recommendation tasks, as they fail to adequately incorporate behavioral information, which intuitively reflects the users’ preferences and is crucial for an effective recommender systems.
-3. PRELIMINARIES
-In this section, we analyze and prove in detail the problems of the content features extracted by pre-trained encoders for recommendation tasks.
-In section 3.1 we introduce an attribution analysis for visual attribution analysis of content features in multimodal recommender systems.
-In section 3.2, we specifically analyze the deficiencies of pre-trained feature encoders.
-In section 3.3 we analyze and prove the problems from a mathematical and theoretical point of view.
-3.1. Attribution Analysis
-Existing methods for analyzing the weights of image features are predominantly based on class activation mapping (CAM) (CAM, ).
-However, the heatmaps generated by CAM only reflect the contribution of certain image regions to the classification task and fail to assess their impact on recommendation models (scoreCAM, ; gScoreCAM, ).
-These limitations restrict our in-depth understanding of modality feature processing in recommender systems.
-Thus we designed the attribution analysis method for analyzing image features in multimodal recommendations based on the characteristics of recommender systems.
-To visually illustrate the issues raised, we introduce a intuitively analysis for the attribution of multimodal features in the recommender system.
-This method aims to explore the effectiveness of content features extracted by pre-trained multimodal feature encoders in multimodal recommender systems.
-It also provides a visual demonstration of the effectiveness of our proposed adapter in bridging the gap between the content features extracted by the pre-trained encoders and the features required by the recommender system.
-To accurately identify which parts of the content features users pay attention to, we use cosine similarity as a measurement.
-By calculating the cosine similarity of each pixel to the behavioral features in the recommendation models, we generate a saliency heatmap.
-This heatmap is created by weighting and summing the masked feature maps based on the calculated similarities.
-This approach visualizes the portions of the content features that are truly effective in the recommendation task, providing a deeper understanding of the role of multimodal features in recommender systems.
-(1) $\text{similarity}(x,y)=\frac{{v_{x}\cdot v_{y}}}{{\|v_{x}\|\cdot\|v_{y}\|}}$
-Similar to the concept behind CAM-based approaches, we choose the target layer for visualization to ensure interpretability (scoreCAM, ; gScoreCAM, ).
-In the case of image encoders, for both ResNet (resnet, ) and ViT (ViT, ) versions, we select the channels closest to the CLIP prediction layer.
-In convolutional neural networks (CNNs), as the layers go deeper, the information represented in the feature maps becomes more abstract and high-level.
-The feature maps of the last ReLU layer capture the most important high-level features in the input image.
-ReLU-activated feature maps contain only positive or zero activations, which helps to emphasize the regions that the model considers more relevant.
-For the ViT series, as the gradients are zero for all layers except for the [CLS] token and only the [CLS] token is used for final predictions in the last layer embedding, we choose the penultimate layer.
-Consequently for the ResNet series, we select the last ReLU layer of the final Bottleneck as the target layer, for the ViT series, we use the second-to-last ResidualAttentionBlock as the target layer.
-In the first stage, we perform a forward pass of the product images through CLIP, obtaining $N$ channels at the target layer denoted as ${A_{i}}^{N}$.
-The purpose is to discriminate the contributions of different targets across different channels.
-Subsequently, for each channel, we upsample it, denoted as $A_{i}^{up}$.
-This is done to generate masked images corresponding to different parts and to calculate the similarity weights of different regions.
-In the second stage, element-wise multiplication is performed between each mask $A_{i}^{up}$ and all color channels $x$ of the input image, and then the result is compared with the corresponding to behavioral features to compute the similarity.
-We name our method Behavior Information CAM (BeCAM) and its saliency map is computed by:
-(2) $M_{\mathrm{BeCAM}}^{i}=\sum_{i}^{N}A_{i}\times\text{similarity}((x\odot A_{i}^{up}),{ID}_{i})$
-where $\odot$ is the Hadamard product.
-3.2. Deficiency Analysis
-Figure 4. An illustration of a theoretical analysis of deficiency analysis
-Figure 5. Comparison of efficient parameter tuning methods. (a) Low-Rank Adaptation (b) Soft Prompt Turning (c) Our proposed adapter.
-By analyzing the generated heatmap (shown in Figure 2), it is clear that the pre-trained feature encoder suffers from the issues of wrong region of interest and insufficient region of interest on the product dataset.
-This can be visualized by the fact that the content features extracted by the pre-trained feature encoders have the problems of information drift and information omission.
-This is a very visible demonstration of the low information value density of the multimedia content itself, resulting the content features are insufficient to reflect users’ preferences.
-Specifically, as shown in Figure 2, although the user wants to buy clothes, the feature encoder incorrectly focuses on the model’s face, resulting in information drift.
-There are also products whose content features miss some key details of the product, such as only focusing on certain details of the clothes and failing to focus on its whole, resulting in information omission.
+Related Work
+Figure 3: Pipeline of the proposed attribution analysis method.
+Multimedia Recommendation
+Collaborative filtering (CF) based methods leverage behavioral similarities for top-k recommendations (Zhang et al.
+2014).
+To improve the performance of CF-based methods, researchers integrate item multimodal content.
+Typically, they employ pre-trained neural networks to extract content features, and then fuse content features with behavioral features to more effectively model user preference.
+(Yuan et al.
+2023).
+For instance, MMGCN (Wei et al.
+2019) creates modality-specific interaction graphs.
+LATTICE (Zhang et al.
+2021) adds links between similar items.
+FREEDOM (Zhou and Shen 2023) constructs graphs using sensitivity edge pruning.
+Recently, self-supervised learning methods like BM3 (Zhou et al.
+2023c) and MGCN (Yu et al.
+2023) further enhance representation quality by reconstructing interaction graphs and adaptively learning content feature importance.
+However, item content information often contains much irrelevant information and exhibiting low value density.
+We argue pre-extracted item content features insufficiently reflect users’ preferences.
+Directly incorporating these features may mislead user preference modeling, resulting in suboptimal recommendation performance.
+Attribution Analysis
+Modality feature attribution analysis methods such as Class Activation Mapping (CAM) (Zhou et al.
+2016), Grad-CAM (Selvaraju et al.
+2017), and Score-CAM (Wang et al.
+2020) generate heatmaps to visualize the decision-making process of convolutional neural networks.
+Grad-CAM uses gradients for better applicability across network architectures.
+Score-CAM weight activation maps with output scores, which eliminates the dependence on gradients.
+However, current attribution analysis methods primarily focus on analyzing the encoder itself and fail to visually analyze items’ content features in recommender systems.
+Most of these methods use gradient information to obtain weights, which reflect more of the encoder focus.
+As a result, they fail to capture the similarity between content features and behavioral features, which directly reflect the user’s preferences.
+Consequently, these methods are unsuitable for use in recommender systems.
+Thus, to better analyze the quality of item content features, we need to design a novel similarity-based attribution analysis method.
+Parameter-Efficient Adaptation
+In multimodal domains, semantic differences limit feature expressiveness, impacting model performance (Khurana et al.
+2023; Rinaldi, Russo, and Tommasino 2023).
+Using pre-trained encoders to extract features in real-world scenarios can result in not accurately reflecting user preferences.
+To address this challenge, researchers have explored techniques such as fine-tuning and adapters to reduce the impact of semantic differences.
+In the context of recommendation scenarios where data is continuously changing and expanding, performing full fine-tuning of the encoder faces huge challenges.
+Due to the above limitations, we consider low-parameter fine-tuning methods to efficiently adapt content features in real time.
+Low-Rank Adaptation (LoRA) (Hu et al.
+2021) integrate low-rank decomposition matrices, while Prompt Tuning (Lester, Al-Rfou, and Constant 2021) uses learnable embedding vectors as hints.
+However, adding adapters to the middle layer of the pre-trained model may lead to information loss.
+Existing methods are not generalized for recommendation tasks.
+They fail to adequately incorporate behavioral information, which is crucial for intuitively reflecting users’ preferences.
+Preliminaries
+Attribution Analysis
+Existing attribution methods are mainly focused on pre-trained encoders, which can only analyze the quality of the content features themselves, fail to reflect their contributions to recommendations.
+Considering that behavioral features directly reflect user preferences in recommendation, we propose a similarity-based attribution analysis.
+This method explores the quality of content features extracted by a pre-trained encoder in recommendations.
+By calculating the cosine similarity of each pixel to the behavioral features, we generate a heatmap that visually identifies the content features relevant to users’ behaviors.
+We follow the core principle of Class Activation Mapping (CAM) to linearly weight feature maps.
+The main difference compared to the previous series of CAM methods is the way in which the linear weights are obtained.
+Our approach consists of two primary stages.
+In the first stage, we take the item image $I$ as input to the corresponding encoder, obtaining $N$ images on the corresponding channels at the target layer denoted as ${a_{l}}$.
+Then we upsample it, denoted as $a_{l}^{up}$.
+This is done to generate masked images corresponding to different parts.
+Considering in all types of encoders, layers closer to the prediction layer contain the highest level and most abstract information (Vaswani et al.
+2017).
+Consequently, for the ResNet family visual encoder, we select the last ReLU layer of the final Bottleneck as the target layer, while for the ViT family visual encoder, we choose the penultimate ResidualAttentionBlock.
+The corresponding content feature $\boldsymbol{x_{l}}$ is extracted from each mask $a_{l}^{up}$.
+Behavioural features $\boldsymbol{h_{i}}$ are obtained from existing multimedia recommendation models.
+The weights of the corresponding content features are obtained by calculating the similarity between them.
+Figure 4: Comparison of parameter tuning methods. (a) Low-Rank Adaptation (b) Soft Prompt Turning (c) BeFA.
+$\text{similarity}(\boldsymbol{x_{l}},\boldsymbol{h_{i}})=\frac{{\boldsymbol{x_{l}}\cdot\boldsymbol{h_{i}}}}{{\|\boldsymbol{x_{l}}\|\cdot\|\boldsymbol{h_{i}}\|}}$ (1)
+Ultimately, the final visualization result is achieved by linearly weighting and summing the feature maps with the similarity. The heatmap $M$ is represented as:
+$M=\sum_{l}^{N}a_{l}\times\text{similarity}(f_{pre}(I_{c}\odot a_{l}^{up}),\boldsymbol{h_{i}})$ (2)
+where $\odot$ is the Hadamard product. $I_{c}$ are the all color channels of the image. $f_{pre}$ denotes the process of extracting features using the corresponding encoder.
+Deficiency Analysis
+From Figure 2 we observed that the pre-trained feature encoder suffers from the issues of wrong region of interest (information drift) and insufficient region of interest (information omission).
+Specifically, although the user wants to buy clothes, the feature encoder incorrectly focuses on the face of the person who wears the cloth, a phenomenon we call information drift.
+Meanwhile, we also found that some of the features did not fully reflect all the characteristics of the item.
+Such as only focusing on certain details of the clothes and failing to focus on its whole, a phenomenon we call information omission.
 Information drift can cause the recommender system to ignore useful and critical information.
 This leads to a loss of important guidance for the recommendation task.
-It also results in a mismatch between recommendations and users’ preferences.
+Which lead a mismatch between recommendations and users’ preferences.
 Information omission can result in important visual or textual cues being ignored or incomplete.
 This causes a loss of critical information needed for the recommendation task.
-These issues collectively compromise the overall effectiveness of the recommender system.
-As a result, directly applying the content features extracted by the pre-trained encoder in the multimodal recommender system will inevitably impair the performance of the recommender system.
-3.3. Theoretical Analysis
-In this section, we reveal the deficiencies of content features extracted by pre-trained feature encoders within the embedding space.
-We analyze the gap between these content features and the ideal features required by the recommendation model.
-Furthermore, we demonstrate that optimizing the consistency between the extracted features and the ideal features necessary for accurate recommendation predictions can significantly enhance recommendation performance.
-This analysis provides theoretical support for the effectiveness of our proposed adapter.
-The embedding space captures the semantic relationships and similarities between features.
-Analyzing features in the embedding space enables a better understanding and interpretation of their relationships.
-Formally, considering a node ${i}$ and its representation $\boldsymbol{e_{i}}$ in the embedding space, which represents the ideal features required by the recommendation model.
-On the basis of the item feature representation $\boldsymbol{e_{i}}$ in the embedding space, the recommender system calculates the probability distribution of the rating or clicking behavior of the user $u$ on the item $i$, which can be viewed as the prediction $P(r_{ui}|\boldsymbol{e_{u}},\boldsymbol{e_{i}})$.
-Where $r_{ui}$ stands for the rating or click behavior of the user $u$ on item $i$, $\boldsymbol{e_{u}}$ represents the user’s feature, and $\boldsymbol{e_{i}}$ signifies the item’s feature.
-Our objective is to find the optimal user feature $e_{u}$ and item feature $\boldsymbol{e_{i}}$ to maximize the posterior distribution $P(r_{ui}|\boldsymbol{e_{u}},\boldsymbol{e_{i}})$.
-In other words, we aim to maximize the posterior distribution for the user feature $\boldsymbol{e_{u}}$ and item feature $\boldsymbol{e_{i}}$ parameters.
-This can be formally described by the following objective function:
-(3) $L(\boldsymbol{e_{u}},\boldsymbol{e_{i}})=\arg\max_{\boldsymbol{x_{u}},\boldsymbol{e_{i}}}P(r_{ui}|\boldsymbol{e_{u}},\boldsymbol{e_{i}})$
-The representation $\boldsymbol{e_{i}^{\prime}}$ extracted by the feature encoder can be seen as the prior distribution.
-However, due to semantic differences between the natural data and e-commerce data, there is inevitably some deviations between the features extracted by the pre-trained feature encoder and the ideal features.
-Let us denote this deviation as $\theta$.
-(4) $\theta=\arccos\left(\frac{\boldsymbol{e_{i}}\cdot\boldsymbol{e_{i}^{\prime}}}{\|\boldsymbol{e_{i}}\|\|\boldsymbol{e_{i}^{\prime}}\|}\right)$
-Analyzing $\theta$ helps understand the deviation between the features extracted by the feature encoder and the ideal features.
-A smaller $\theta$ indicates that the two representations are very close in the embedding space.
-This implies a high consistency between the features extracted by the feature encoder and the ideal features, which benefits the recommendation system in accurately capturing user interests and item similarities.
-By reducing $\theta$, the features become closer to the ideal features, thereby enhancing the performance of the recommendation system.
-The part $p$ of the representation $\boldsymbol{e_{i}^{\prime}}$ extracted by the encoder that is truly effective for the recommendation task can be represented as follows:
-(5) $p=\left(\boldsymbol{e_{i}^{\prime}}\cdot\frac{\boldsymbol{e_{i}}}{\|\boldsymbol{e_{i}}\|}\right)\frac{\boldsymbol{e_{i}}}{\|\boldsymbol{e_{i}}\|}=\boldsymbol{e_{i}^{\prime}}\cdot\cos(\theta)$
-When the deviation $\theta$ is large, as $\boldsymbol{e_{i}^{\prime}}$ in Figure 4, content features exhibit the issues of information omission, causing the effective length of $p$ deviates significantly from $\boldsymbol{e_{i}}$.
-This results in a lack of crucial information in the extracted features, reflected as insufficient region of interest in the heatmap.
-Conversely, when $\theta$ is too large,as $\boldsymbol{e_{j}}$ in Figure 4, content features exhibit the issues of information drift, with $p$ being located in the wrong quadrant and its effective direction opposing $\boldsymbol{e_{i}}$.
-This introduces a large amount of incorrect information, reflected as incorrect region of interest in the heatmap.
-In such circumstances, the recommendation system is easily to make incorrect recommendations, because the feature representation contradicts the users’ preferences, failing to appropriately reflect the characteristics of items or users.
-In the recommendation task, we aim for a smaller deviation between $\boldsymbol{e_{i}^{\prime}}$ and $\boldsymbol{e_{i}}$.
-To formalize this objective, we define an error function $f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})$ to measure the gap between the representations:
-(6) $f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})=\left(1-\frac{{\boldsymbol{e_{i}^{\prime}}\cdot\boldsymbol{e_{i}}}}{{\|\boldsymbol{e_{i}^{\prime}}\|\|\boldsymbol{e_{i}}\|}}\right)$
-We seek to minimize the expected deviation measure $\Delta$. We can define $\Delta$ as:
-(7) $\Delta=\mathbb{E}_{P(\boldsymbol{e_{i}^{\prime}})}[f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})]$
-The function $f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})$ measures the deviation between representations.
-In recommendation tasks, the quality of the feature representation $\boldsymbol{e_{i}^{\prime}}$ directly affects the accuracy and effectiveness of recommendations.
-A decline in the quality of $\boldsymbol{e_{i}^{\prime}}$, indicated by an increase in the value of the error function $f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})$, donates an augmented diversity between the representation extracted by the pre-trained encoder $\boldsymbol{e_{i}^{\prime}}$ and the ideal representation $\boldsymbol{e_{i}}$.
-Consequently, this leads to an increase in the expected deviation $\Delta$, subsequently influencing the posterior distribution $P(r_{ui}|\boldsymbol{e_{u},e_{i}})$.
-Such circumstances may impair the recommendation system’s ability to accurately capturing the users’ preferences, thereby reducing the accuracy and efficacy of recommendations.
-By minimizing $\Delta$, we can ensure that the representation $\boldsymbol{e_{i}^{\prime}}$ are closer to the ideal representation $\boldsymbol{e_{i}}$ in expectation.
-This adaptation of content features enhances the consistency with the ideal representation and thus improves the performance of the recommender system.
-Figure 6.
-Visual analysis of the effect of adapter on feature purification.
-(a) is the original image.
-(b) is the similarity heatmap of the visual features extracted by the pre-trained CLIP.
-(c) (d) (e) are the similarity heatmaps of the visual features adapted by BeFA applied on FREEDOM, BM3 and LATTICE respectively
-Table 1. Statistics of the experimental datasets
-Dataset #User #Item #Behavior Density
-TMALL 13,104 7,848 151,928 0.148%
-Microlens 46,420 14,079 332,730 0.051%
-H&M 43,543 16,915 369,945 0.050%
-4. feature adapter
-To address the discovered issues, we proposed Behavior-driven Feature Adapter (BeFA).
-This is a adapter for adapting multimodal content features that are employed by multimodal recommender systems.
-Specifically, the content features are first extracted by a pre-trained encoder, and then the content features are adapted by BeFA, and then fed into the multimodal recommender system for item and user modeling.
-BeFA and the downstream recommender system share the optimization objective and use End-to-End training.
-4.1. Problem Formulation
-Let $u\in\mathcal{U}$ and $i\in\mathcal{I}$ denote the user and item, respectively.
-The input behavioral features for user $u$ and item $i$ are represented as $\mathrm{E_{id}}\in\mathbb{R}^{d\times(|U|+|I|)}$, where $d$ is the embedding dimension.
-Each item modality feature is denoted as $\mathrm{E_{i,m}}\in\mathbb{R}^{d_{m}\times|I|}$, where $d_{m}$ is the dimension of the features, $m\in\mathcal{M}$ represents the modality, and $\mathcal{M}$ is the set of modalities.
-In this paper, we primarily focus on visual and textual modalities, denoted by =$\mathcal{M}=\{v,t\}$.
-It’s worth noting that our approach is adaptable to incorporate more than two modalities.
-Next, the users’ historical behavior feature is represented as $R\in\{0,1\}^{|U|\times|I|}$, where each entry $R_{u,i}=1$ if user $u$ clicked item $i$, otherwise $R_{u,i}=0$ This historical interaction data $R$ can naturally be interpreted as a user-item bipartite graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$.
-In $\mathcal{G}$ , the vertex set $\mathcal{V}=\mathcal{U}\cup\mathcal{I}$ donates all items and users.
-The edge set =$\mathcal{E}=\{(u,i)|u\in\mathcal{U},i\in\mathcal{I},R_{u,i}=1\}$ represents the interacted users and items.
-The objective of multimedia recommendation is to accurately predict users’ preferences by ranking items for each user based on predicted preference scores $\hat{y}_{u,i}$.
-Table 2. Performance Comparison on Different Recommender Models. The t-tests validate the significance of performance improvements with p-value $\leq$ 0.05.
+These issues compromise the overall performance of the recommender system.
+As a result, directly applying the pre-extracted content features will inevitably impair the performance of the recommender system.
+In the Appendix, we perform a theoretical analysis to further illustrate the point.
+Feature Adapter
+To address the discovered issues, we propose Behavior-driven Feature Adapter (BeFA).
+The purpose of this adapter is to efficiently adapt content features to improve their quality in recommendations, thus improving the performance of recommendations.
+Specifically, the content features are first extracted by a pre-trained encoder and then adapted by BeFA.
+Then it is fed into the multimodal recommender system for item and user modeling.
+BeFA and the downstream recommender system share the optimization objective and adopt End-to-End training strategy.
+Problem Formulation
 TMALL Microlens H&M
@@
 BM3+BeFA 0.0212 0.0319 0.0115 0.0144 0.0566 0.0911 0.0314 0.0412 0.0266 0.0391 0.0157 0.0190
-Imporve 12.17% 7.05% 12.75% 9.09% 10.98% 7.05% 12.95% 9.87% 30.39% 22.19% 37.72% 31.94%
+Improve 12.17% 7.05% 12.75% 9.09% 10.98% 7.05% 12.95% 9.87% 30.39% 22.19% 37.72% 31.94%
 LATTICE 0.0238 0.0356 0.0134 0.0167 0.0553 0.0886 0.0308 0.0402 0.0289 0.0427 0.0161 0.0197
 LATTICE+BeFA 0.0260 0.0403 0.0183 0.0183 0.0593 0.0943 0.0328 0.0427 0.0317 0.0498 0.0171 0.0217
-Imporve 9.24% 13.20% 36.57% 9.58% 7.23% 6.43% 6.49% 6.22% 9.69% 16.63% 6.21% 10.15%
+Improve 9.24% 13.20% 36.57% 9.58% 7.23% 6.43% 6.49% 6.22% 9.69% 16.63% 6.21% 10.15%
 FREEDOM 0.0212 0.0340 0.0113 0.0148 0.0474 0.0774 0.0262 0.0348 0.0348 0.0526 0.0188 0.0234
@@
 Improve 4.82% 3.95% 5.19% 4.68% 1.94% 2.88% 2.63% 3.17% 10.35% 8.20% 10.29% 9.16%
-CLIP Avg Imporve 11.39% 8.62% 18.71% 9.55% 6.57% 5.38% 7.14% 6.25% 16.99% 14.46% 18.61% 16.77%
+CLIP Avg Improve 11.39% 8.62% 18.71% 9.55% 6.57% 5.38% 7.14% 6.25% 16.99% 14.46% 18.61% 16.77%
 BM3 0.0184 0.0299 0.0097 0.0129 0.0508 0.0842 0.0279 0.0373 0.0195 0.0304 0.0107 0.0135
 BM3+BeFA 0.0224 0.0322 0.0125 0.0152 0.0537 0.0877 0.0299 0.0395 0.0248 0.0378 0.0149 0.0183
-Imporve 21.74% 7.69% 28.87% 17.83% 5.71% 4.16% 7.17% 5.90% 27.18% 24.34% 39.25% 35.56%
+Improve 21.74% 7.69% 28.87% 17.83% 5.71% 4.16% 7.17% 5.90% 27.18% 24.34% 39.25% 35.56%
 LATTICE 0.0252 0.0374 0.0139 0.0173 0.0580 0.0953 0.0320 0.0426 0.0293 0.0439 0.0164 0.0202
 LATTICE+BeFA 0.0266 0.0403 0.0147 0.0184 0.0633 0.1021 0.0340 0.0451 0.0316 0.0498 0.0171 0.0218
-Imporve 5.56% 7.75% 5.76% 6.36% 9.14% 7.14% 6.25% 5.87% 7.85% 13.44% 4.27% 7.92%
+Improve 5.56% 7.75% 5.76% 6.36% 9.14% 7.14% 6.25% 5.87% 7.85% 13.44% 4.27% 7.92%
 FREEDOM 0.0197 0.0319 0.0107 0.0140 0.0613 0.0976 0.0337 0.0440 0.0364 0.0553 0.0192 0.0241
@@
 Improve 3.38% 2.73% 5.56% 1.65% 1.30% 0.93% 0.26% 0.40% 11.36% 6.85% 12.84% 9.93%
-ImageBind Avg Imporve 15.54% 9.09% 17.99% 12.53% 5.18% 3.77% 4.83% 4.12% 15.24% 13.87% 19.56% 17.92%
-4.2. Behavior-driven Feature Adapter
-Due to the shortcomings of pre-trained feature encoders, their extracted content features contain a significant amount of irrelevant and erroneous information.
+ImageBind Avg Improve 15.54% 9.09% 17.99% 12.53% 5.18% 3.77% 4.83% 4.12% 15.24% 13.87% 19.56% 17.92%
+Table 1: Performance Comparison on Different Recommender Models. The $t$-tests validate the significance of performance improvements with $p$-value $\leq$ 0.05.
+Let $u\in\mathcal{U}$ and $i\in\mathcal{I}$ denote the user and item, respectively.
+The input behavioral features for user $u$ and item $i$ are represented as $\mathrm{E_{b}}\in\mathbb{R}^{d\times(|\mathcal{U}|+|\mathcal{I}|)}$, where $d$ is the embedding dimension.
+Each row represents the user embedding $\boldsymbol{h_{u}}$ and item embedding $\boldsymbol{h_{i}}$.
+They are initialised with ID information (He and McAuley 2016).
+Each item content feature is denoted as $\boldsymbol{e_{i,m}}\in\mathbb{R}^{d_{m}}$, where $d_{m}$ is the dimension of the features.
+$m\in\mathcal{M}$ represents the modality and $\mathcal{M}$ is the set of modalities.
+Multimedia recommendation aims to rank items for each user by predicting preference scores $\hat{y}_{u,i}$, which indicates the possibility of interaction between the user $u$ and the item $i$.
+The recommended process can be viewed as a function $f(\cdot)$.
+This process can be viewed simply as:
+$\hat{y}_{u,i}={f}(\boldsymbol{h_{u}},\boldsymbol{h_{i}},\boldsymbol{e_{i,m}})$ (3)
+Existing multimodal recommendation methods typically use pre-trained encoders to extract content features $\boldsymbol{e_{i,m}}$.
+Due to the deficiencies of the pre-trained encoder, a large deviation in the content features $\boldsymbol{e_{i,m}}$ leads to a bias in the prediction score $\hat{y}_{u,i}$, which affects the recommendation results.
+Therefore, we designed BeFA to adapt the content features to make them more suitable for recommendation.
+BeFA improves the recommendation results by adjusting $\boldsymbol{e_{i,m}}$, which can be viewed as a function $g(\cdot)$.
+The adjusted predicted preference scores is donated:
+$\hat{y}_{u,i}={f}[\boldsymbol{h_{u}},\boldsymbol{h_{i}},g(\boldsymbol{h_{i}},\boldsymbol{e_{i,m}})]$ (4)
+Behavior-driven Feature Adapter
+Due to the shortcomings of pre-trained feature encoders, their extracted content features contain amount of irrelevant and erroneous information.
 To better utilize modality information,we propose Behaviour-driven Feature Adapter(BeFA) for adapting content features.
-Firstly, we decouple the original item content features $\mathrm{E_{i,m}}$ into features in the decoupled feature space $\mathrm{\dot{E}_{i,m}}$:
-(8) $\dot{\mathrm{E}}_{i,m}=\mathrm{W}_{1}\mathrm{E}_{i,m}+\mathrm{b}_{1},$
-where $\mathbf{W}_{1}\in\mathbb{R}^{d\times d_{\lambda}}$ and $\mathbf{b}_{1}\in\mathbb{R}^{d_{m}}$ represent trainable transformation matrix and bias vector, respectively.
-Considering that the behavioral information fully reflects the users’ preference, we filter the preference-related content features with the help of behavioral information guidance:
-(9) $\ddot{\mathrm{E}}_{i,m}=f_{gate}^{m}(\mathrm{E}_{i,id},\dot{\mathrm{E}}_{i,m})=\mathrm{E}_{i,id}\odot\sigma(\mathrm{W}_{2}\dot{\mathrm{E}}_{i,m}+\mathrm{b}_{2}),$
-where $\mathbf{W}_{2}\in\mathbb{R}^{d_{\lambda}\times d_{\lambda}}$ and $\mathbf{b}_{2}\in\mathbb{R}^{d_{m}}$ are trainable parameters.
-Here, $\odot$ denotes element-wise multiplication, and $\sigma$ is the Tanh non-linear transformation.
-Finally, the decoupled content features are recombined.
-Given that behavioral information reflects user preferences, we determine the weights of different decoupled content features based on the guidance provided by this behavioral information:
-(10) $\overline{\mathrm{E}}_{i,m}=f_{merge}^{m}(\mathrm{E}_{i,id},\ddot{\mathrm{E}}_{i,m})=\mathrm{E}_{i,id}\odot\sigma(\mathrm{W}_{3}\ddot{\mathrm{E}}_{i,m}+\mathrm{b}_{3}),$
-where $\mathbf{W}_{3}\in\mathbb{R}^{d_{\lambda}\times d}$ and $\mathbf{b}_{3}\in\mathbb{R}^{d}$ are trainable parameters.
-Similarly, $\odot$ denotes element-wise multiplication, and $\sigma$ is the sigmoid non-linear transformation.
-It is important to note that although Equations 9 and 10 are formally similar, they achieve different effects.
+Firstly, we decouple the original item $i$ content features $\boldsymbol{e_{i,m}}$ in the decoupled feature space:
+$\dot{\boldsymbol{e}}_{i,m}=\mathbf{W}_{1}\boldsymbol{e}_{i,m}+\mathbf{b}_{1},$ (5)
+where $\mathbf{}{W}_{1}\in\mathbf{}{R}^{d_{m}\times d_{a}}$ and $\mathbf{b}_{1}\in\mathbb{R}^{d_{a}}$ represent trainable transformation matrix and bias vector.
+$d_{a}$ is a hyper-parameter represents the dimension of the decoupled space.
+Considering that the behavioral information reflects the users’ preference (Yuan et al.
+2023), we filter the preference-related content features with the guidance of behavioral information.
+The filter function $f_{gate}^{m}(\cdot)$ is represented as:
+$\ddot{\boldsymbol{e}}_{i,m}=f_{gate}^{m}(\boldsymbol{b}_{i},\dot{\boldsymbol{e}}_{i,m})=\boldsymbol{b}_{i}\odot\tanh(\mathbf{W}_{2}\dot{\boldsymbol{e}}_{i,m}+\mathbf{b}_{2}),$ (6)
+where $\mathbf{W}_{2}\in\mathbb{R}^{d_{a}\times d_{a}}$ and $\mathbf{b}_{2}\in\mathbb{R}^{d_{a}}$ are trainable parameters, $\boldsymbol{b_{i}}$ denotes the behavioral information of item i.
+Here $\tanh(\cdot)$ is the Tanh non-linear transformation.
+Finally, the decoupled content features are selectively recombined.
+By combining different decoupled features with behavioral guidance, we aim to enhance the system’s ability to capture nuanced and contextually relevant aspects of the content.
+The combination function $f_{merge}^{m}$ is represented as:
+$\overline{\boldsymbol{e}}_{i,m}=f_{merge}^{m}(\boldsymbol{b}_{i},\ddot{\boldsymbol{e}}_{i,m})=\boldsymbol{b}_{i}\odot\sigma(\mathbf{W}_{3}\ddot{\boldsymbol{e}}_{i,m}+\mathbf{b}_{3}),$ (7)
+where $\mathbf{W}_{3}\in\mathbb{R}^{d_{a}\times d_{m}}$ and $\mathbf{b}_{3}\in\mathbb{R}^{d_{m}}$ are trainable parameters.
+$\sigma(\cdot)$ is the sigmoid non-linear transformation.
+It is important to note that although Equations 6 and 7 are formally similar, they achieve different effects.
 Moreover, we introduce ReLU activation functions and Dropout between each transformation layer.
@@
 Additionally, it enhances the model’s generalization, making it more suitable for practical application scenarios.
-5. EXPERIMENTS
-In this section, we conduct extensive experiments to evaluate the performance of the proposed adapter on three public datasets to answer following research questions.
-  • RQ1: How effective is BeFA applied to existing multimodal recommender systems?
-  • RQ2: How effective is BeFA in adapting content features?
-  • RQ3: Why purifying modality information can achieve better recommendation performance?
-  • RQ4: How do hyper-parameter settings impact the performance of BeFA?
-5.1. Experimental Settings
-5.1.1. Datasets
+Experiments
+Experimental Settings
+Datasets
 We conducted experiments on three publicly available datasets: (a) TMALL [^1] https://tianchi.aliyun.com/dataset/140281; (b) Microlens [^2] https://recsys.westlake.edu.cn/MicroLens-50k-Dataset/; and (c) H&M [^3] https://www.kaggle.com/datasets/odins0n/handm-dataset-128x128.
-We performed 10-core filtering on the raw data.
-The detailed information of the filtered data is presented in Table 1.
-For multimodal information, we utilized pre-trained CLIP (clip, ) and ImageBind (imagebind, ) to generate aligned multimodal features.
-5.1.2. Comparative Model Evaluation
-To verify the prevalence of feature encoder defects, we used ViT-B/32-based CLIP (clip, ) and ImageBind (imagebind, ) for our experiments.
-CLIP is a general-purpose feature encoder capable of handling a wide range of data types and application domains.
-It realizes cross-modality feature representations and exhibits highly efficient performance, making it widely adopted in practical applications and performing well across various tasks.
-ImageBind is the latest cross-modality feature encoder, which can handle a wider variety of modality information, and achieves a new SOTA performance in the emerging zero-shot recognition task for various modalities, even outperforming the previous specialized models trained specifically for these modalities.
-To evaluate the effectiveness of our proposed adapter, we applied it to several representative multimodal recommendation models, including LATTICE (lattice, ), BM3 (bm3, ), FREEDOM (FREEDOM, ), and MGCN (mgcn, ) for comparison.
-Additionally, we compared our adapter with existing efficient parameter tuning methods such as LoRA (lora, ) and Soft-Prompt Tuning (softprompt, ).
-Table 3.
-Performance Comparison with other Efficient Parameter Adaptation methods.
+The detailed information is presented in the Appendix.
+For multimodal information, we utilized pre-trained CLIP (Radford et al.
+2021) and ImageBind (Girdhar et al.
+2023) to extract aligned multimodal features.
+Comparative Model Evaluation
+To verify the prevalence of feature encoder defects, we used ViT-B/32-based CLIP and ImageBind for our experiments.
+CLIP and ImageBind are advanced cross-modality feature encoders.
+CLIP known for its broad applicability and efficiency, and ImageBind achieving SOTA performance in zero-shot recognition across various modalities.
+To evaluate the effectiveness of BeFA, we applied it to several representative multimodal recommendation models, including LATTICE (Zhang et al.
+2021), BM3 (Zhou et al.
+2023c), FREEDOM (Zhou and Shen 2023), and MGCN (Yu et al.
+2023).
+Additionally, we compared BeFA with existing efficient parameter tuning methods, including LoRA (Hu et al.
+2021) and Soft-Prompt Tuning (Lester, Al-Rfou, and Constant 2021).
+Evaluation Protocols & Implementation
+Due to the length of the article, the details are shown in the Appendix.
+Overall performance
+Figure 5: Visualization analysis of the effect of adapter on feature purification. The shade of the colour represents the amount of attention weight.
+TMALL Microlens H&M
+Dadasets R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20
+BM3 0.0189 0.0298 0.0102 0.0132 0.0510 0.0851 0.0278 0.0375 0.0204 0.0320 0.0114 0.0144
+BM3+LoRA 0.0215 0.0333 0.0117 0.0150 0.0510 0.0850 0.0280 0.0376 0.0191 0.0290 0.0110 0.0135
+BM3+SoftPrompt 0.0193 0.0298 0.0102 0.0131 0.0517 0.0861 0.0285 0.0382 0.0202 0.0311 0.0114 0.0142
+BM3+BeFA 0.0212 0.0319 0.0115 0.0144 0.0566 0.0911 0.0314 0.0412 0.0266 0.0391 0.0157 0.0190
+LATTICE 0.0238 0.0356 0.0134 0.0167 0.0553 0.0886 0.0308 0.0402 0.0289 0.0427 0.0161 0.0197
+LATTICE+LoRA 0.0254 0.0384 0.0147 0.0183 0.0550 0.0884 0.0304 0.0399 0.0289 0.0434 0.0163 0.0201
+LATTICE+SoftPrompt 0.0266 0.0389 0.0148 0.0182 0.0539 0.0889 0.0294 0.0393 0.0301 0.0459 0.0166 0.0207
+LATTICE+BeFA 0.0260 0.0403 0.0183 0.0183 0.0593 0.0943 0.0328 0.0427 0.0317 0.0498 0.0171 0.0217
+FREEDOM 0.0212 0.0340 0.0113 0.0148 0.0474 0.0774 0.0262 0.0348 0.0348 0.0526 0.0188 0.0234
+FREEDOM+LoRA 0.0197 0.0323 0.0106 0.0141 0.0476 0.0774 0.0264 0.0349 0.0352 0.0533 0.0190 0.0237
+FREEDOM+SoftPrompt 0.0215 0.0342 0.0118 0.0154 0.0465 0.0769 0.0258 0.0345 0.0406 0.0577 0.0224 0.0268
+FREEDOM+BeFA 0.0243 0.0364 0.0131 0.0164 0.0503 0.0814 0.0279 0.0368 0.0409 0.0583 0.0226 0.0271
+MGCN 0.0249 0.0380 0.0135 0.0171 0.0618 0.0972 0.0342 0.0442 0.0367 0.0549 0.0204 0.0251
+MGCN+LoRA 0.0260 0.0391 0.0141 0.0179 0.0598 0.0963 0.0335 0.0438 0.0367 0.0554 0.0203 0.0252
+MGCN+SoftPrompt 0.0260 0.0391 0.0144 0.0180 0.0597 0.0955 0.0334 0.0435 0.0375 0.0557 0.0207 0.0254
+MGCN+BeFA 0.0261 0.0395 0.0142 0.0179 0.0630 0.1000 0.0351 0.0456 0.0405 0.0594 0.0225 0.0274
+Avg Improve 2.44% 1.71% 7.89% 0.48% 6.08% 4.98% 6.25% 5.67% 11.11% 9.59% 12.58% 11.44%
+Table 2: Performance Comparison with other Efficient Parameter Adaptation methods.
 The best result is in boldface and the second best is underlined.
 The $t$-tests validate the significance of performance improvements with $p$-value $\leq$ 0.05.
-TMALL Microlens H&M
-Dadasets R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20
-[…]
-BM3+LoRA 0.0215 0.0333 0.0117 0.0150 0.0510 0.0850 0.0280 0.0376 0.0191 0.0290 0.0110 0.0135
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-FREEDOM 0.0212 0.0340 0.0113 0.0148 0.0474 0.0774 0.0262 0.0348 0.0348 0.0526 0.0188 0.0234
-[…]
-[…]
-FREEDOM+BeFA 0.0243 0.0364 0.0131 0.0164 0.0503 0.0814 0.0279 0.0368 0.0409 0.0583 0.0226 0.0271
-[…]
-MGCN+LoRA 0.0260 0.0391 0.0141 0.0179 0.0598 0.0963 0.0335 0.0438 0.0367 0.0554 0.0203 0.0252
-[…]
-[…]
-Avg Improve 2.44% 1.71% 7.89% 0.48% 6.08% 4.98% 6.25% 5.67% 11.11% 9.59% 12.58% 11.44%
-i) Multimodal Recommendation Models:
-  • LATTICE (lattice, ): This method designs a new modality-aware structure learning layer that learns the item-item structure of each modality and aggregates multiple modalities to obtain latent item graphs.
-Based on the learned latent graphs, graph convolution is performed to explicitly inject higher-order item affinities into item representations.
-  • BM3 (bm3, ): This method proposes a simple self-supervised learning framework that generates target views of users or items for comparison learning without negative samples.
-  • FREEDOM (FREEDOM, ): This method devises a degree-sensitive edge pruning method to denoise user-item interaction graphs, rejecting potentially noisy edges with high probability when sampling the graph.
-  • MGCN (mgcn, ): This method introduces a behavior-aware fuser to effectively model user preferences by dynamically learning the relative significance of various modality features.
-Additionally, it integrates the fuser with a self-supervised auxiliary task aimed at maximizing the mutual information between the fused multimodal features and behavioral features.
-This approach enables the concurrent capture of complementary and supplementary preference information.
-ii) Efficient Parameter Tuning Methods
-  • Low-Rank Adaptation (LoRA) (lora, ): This method aims to enhance the representation of a model while keeping the number of model parameters relatively small.
-It provides additional learning power by introducing lower rank matrices to linearly transform the hidden representation of the original model.
-This approach strikes a balance between parameter efficiency and model performance.
-  • Soft Prompt Turning (softprompt, ): This method enhances model performance by fine-tuning small, task-specific embeddings (prompts) inserted into the input, thereby adapting pre-trained language models to new tasks without altering their original weights.
-5.1.3. Evaluation Protocols
-For a fair comparison, we follow the evaluation settings in (lattice, ; bm3, ) with the same 8:1:1 data splitting strategy for training, validation and testing.
-Besides, we follow the all-ranking protocol to evaluate the top-K recommendation performance and report the average metrics for all users in the test set: $\mathrm{R@K}$ and $\mathrm{N@K}$, which are abbreviations for $\mathrm{Recall@K}$ (recall, ) and $\mathrm{NDCG@k}$ (NDCG, ), respectively.
-5.1.4. Implementation Details
-We implement MMRec [^4] https://github.com/enoche/MMRec (mmrec, ) based on PyTorch, which is a unified public repository designed for multimodal recommendation methods.
-To ensure fair comparison, we employed the Adam optimizer to optimize all methods and referred to the best hyperparameter settings reported in the original baseline paper.
-For general settings, we initialized embeddings with Xavier initialization of dimension 64, set the regularization coefficient to $\lambda_{E}=10^{-4}$, and batch size to $B=2048$.
-Early stopping and total epochs are fixed at 10 and 1000, respectively.
-We selecte the best model with the highest $\mathrm{Recall@20}$ metric on the validation set and reported metrics on the test set accordingly.
-5.2. Overall performance(RQ1)
-5.2.1. Effectiveness
-(a) Comparison of model performance before and after applying BeFA
-Table 2 shows the performance comparison of BeFA applied to existing multimodal recommendation models on three datasets. From the table,we find several observations:
-  • Our adapter significantly improved recommendation performance on all three datasets.
-Specifically, our adapter has achieved an average improvement of 9.07% and 11.02% over the baseline in terms of $\mathrm{Recall@20}$ and $\mathrm{NDCG@20}$.
-This suggests that the content features extracted by the pre-trained encoder do have deficiencies that affect recommendation performance.
-Our processing of content features is very efficient and effectively reduces modality noise.
+The effectiveness of BeFA.
+Table 1 demonstrates that our adapter enhances recommendation performance across all three datasets.
+Specifically, our adapter has achieved average improvements of 9.07% and 11.02% over the baseline in terms of Recall@20 and NDCG@20.
+This suggests that the pre-extracted content features do have deficiencies that negatively impact recommendation performance.
+Our processing of content features effectively reduces modality noise.
 This helps the multimodal recommendation models to better utilize the modality information and thus improve the recommendation performance.
-  • Our adapter has significant performance improvements for all encoders.
-All pre-trained feature encoders are generally defective.
-We use CLIP and ImageBind as an example, and applying our adapter to the multimodal features extracted by these pre-trained feature encoders has led to a significant improvement in the performance of the recommender system.
-This suggests that the deficiencies are mainly due to the pre-training data, indicating that the content features extracted by the pre-trained encoders generally do not reflect the users’ preferences.
-  • More powerful encoder can extract more sufficient content information.
-ImageBind has better overall preformance than CLIP when applied to recommendation models.
-This indicates that more advanced encoders are able to extract more sufficient content information.
-Thus better modelling of users and items to improve the performance of recommendations.
-(b) Performance comparison with existing parameter tuning methods
-Table 4. The training cost. #Param: number of tunable parameters, Time/E: averaged training time for one epoch, ’s’ means seconds. The results are collected on a GeForce RTX 4090 GPU.
+Furthermore, Table 2 shows the performance of our adapter is better than existing efficient parameter tuning methods, highlighting its superior suitability for recommendation tasks.
+This suggests that our adapter uses behavioral information to better capture user-preferred content features, thus improving the performance.
+The generalization of BeFA.
+From Table 1, we observe that BeFA improves performance across various multimodal recommendation models, highlighting its effectiveness as a generic plugin for all multimodal recommendations.
+Existing multimedia recommendation methods ignore this problem, thus leading to suboptimal results.
+Enhancing content features by BeFA, can make the content features better reflect the users’ preference information.
+Thus make the existing methods more accurately model the users’ preference.
+Meanwhile, BeFA enhances performance for all types of encoders, indicating thar pre-trained encoders are generally defective.
+It can also be found that although more advanced feature encoder (ImageBind) can be used to extract information more comprehensively.
+The extracted features still contain a large amount of irrelevant information, resulting in generally do not accurately reflect the user’s preference.
 Dataset Method #Param. Time/E
@@
 BM3+BeFA +0.20M +0.12s
-Figure 7.
-The effect of LoRA and SoftPrompt Turning on the adaptation of features.
-(a) is the original heatmap, (b) is the heatmap of visual features adapted by LoRA, (c) is the heatmap of visual features adapted by SoftPrompt Turning and (d) is the heatmap of visual features adapted by BeFA.
-Table 3 shows the comparison between our adapter and existing efficient parameter adaptation methods. From the table,we find several observations:
-  • The performance of our adapter is overall significantly better than existing efficient parameter tuning methods.
-This indicates that our adapter is better suited for the recommendation task than generalized efficient parameter tuning methods.
-It suggests that our adapter employs behavioral information guidance to more effectively obtain content features related to users’ preferences, thereby improving the effectiveness of recommendations.
-  • Existing parameter tuning approaches are not generalized for recommendation tasks.
-The performance decreases in some cases instead, which indicates that these existing generalized parameter tuning methods are not necessarily suitable for the recommendation tasks.
-This highlights the importance of incorporating behavioral information guidance in recommendation tasks to improve model performance.
-Figure 8.
-The gaps in the effectiveness of our adapter on different datasets.
-(a) is the original image.
-(b) is the heatmap of the visual features extracted by the pre-trained CLIP.
-(c) is the heatmap of visual features adapted by BeFA.
-[…]
-HM_Microlens
-Figure 9. The Distribution of representations in different multimodal recommendation model in visual modalities before and after adaptation by BeFA.
-5.2.2. Efficiency
-We also analyze the efficiency of our adapter with other existing efficient parameter tuning methods on the BM3 model.
-The specific model parameter counts and training costs are shown in Table 4.
-Compared to the overall model training time, our training time increases by only about 15%, indicating that our method is not only effective but also efficient.
-Although the design of our adapter is more complex compared to LoRA and Soft-Prompt Turning the increase in the number of parameters after using our adapter is still small compared to the overall parameters of the recommender system itself, ranging from 0.93% to 2.09%.
-In terms of training costs, the increase in overall training time after using our adapter is still within an acceptable range.
-The training time for the recommender system using our adapter is comparable to that of using LoRA and Soft-Prompt Tuning.
-This suggests that although our adapter introduces additional parameters, it does not impose a significant burden on the training process.
-The substantial improvement in performance of our adapter highlights its efficiency, trading a small number of parameters and training time gains for a significant improvement in recommendation accuracy.
-5.3. Visualization Analysis(RQ2&RQ3)
-5.3.1. Case Study
-(a)Effectiveness with different models
-To visually demonstrate the effect of our adapter on content features, we applied it to the proposed visualization attribution analysis.
-Specifically, before calculating the similarity between content features and behavioral features in the attribution analysis, we processed the content features with our adapter.
-We then computed similarity weights with the behavioral features and generated new heatmaps using these newly calculated weights with the corresponding masked images.
-As shown in Figure 6, by analysing the heatmap before and after our adapter, we can clearly find that our adapter’s enhancement of content features is very obvious.
-Our adapter significantly improves the attention areas of features extracted by the pre-trained feature encoder, addressing issues of information drift and information omission.
-The adapted features focus more on the recommended products, with a marked decrease in attention to background elements unrelated to the recommendation.
-This reduction in noise enhances feature representation and consequently improves recommendation performance.
-Additionally, the adapted features more comprehensively capture product details, significantly increasing the coverage of the area of interest for products.
-This comprehensive extraction of product details not only improves feature recognition but also enhances the recommender system’s ability to model product features.
-This means that the recommender system is able to more accurately capture the core features of the product, thus providing more relevant and accurate results when making recommendations.
-Overall,our adapter enables multimodal recommender systems to better utilize content features by using behavioral information to guide feature adaptation, thus significantly improving recommendation performance.
-(b) Effectiveness with different encoders
+Table 3: The training cost. #Param: number of tunable parameters, Time/E: averaged training time for one epoch.
+The efficiency of BeFA.
+From Table 3, we observe that our adapter is more complex compared to LoRA and Soft-Prompt Tuning.
+However, the increase in the number of parameters after using our adapter remains small, ranging from 0.93% to 2.09% of the overall recommender system’s parameters.
+The increase in overall training time is only about 15%, which is comparable with other existing methods.
+This suggests that although our adapter is more complex and introduces additional parameters, it does not impose a burden on the training process.The substantial improvement in performance of our adapter highlights its efficiency.
+It trades a small increase in parameters and training time for a improvement in recommendation performance.
+Visualization Analysis
+Figure 6: Comparison between different parameter-efficient adaptation methods.
+BeFA guides content features to more accurately reflect user preferences.
+To visually demonstrate the effect of our adapter on content features, we applied the purified features to the proposed visualization attribution analysis.
+As shown in Figure 5, our adapter addresses issues of information drift and information omission.
+The adapted features focus more on the recommended items, with a marked decrease in attention to background elements unrelated to the recommendation.
+Additionally, the adapted features more comprehensively capture item details.
+This comprehensive extraction of item details improves feature recognition, which enhances the recommender system’s ability to model item features.
+Thus provides more relevant and accurate results when making recommendations.
+We also visualized the effect of BeFA on the distribution of visual representations.
+We find that the adapted features more accurately reflect the detailed information of the items, and the discriminability of the features is clearly improved.
+This is reflected in the distribution as a noticeable improvement in homogeneity (Wang et al.
+2022).
+BeFA is generalized with different encoders.
 We utilize the content features extracted by two representative multimodal feature encoders, the ViT-32/B version of CLIP and ImageBind, and analyze the heatmaps of the content features within different downstream recommendation models after applying BeFA.
 Our analysis reveals that the content features extracted by both encoders exhibit issues of information drift and information omission.
-After applying BeFA, these content features are better focused on the recommended products themselves.
+After applying BeFA, these content features are better reflect the recommended items themselves.
 This demonstrates that the deficiencies of pre-trained feature encoders are common in multimodal recommendation tasks, while our adapter effectively adapts to the content features extracted by different feature encoders, highlighting its general applicability.
-(c) Comparison with existing adaptation methods
+Figure 7: The effectiveness of BeFA on different scenarios. The top image is the heatmap and the bottom one is the heatmap after applying BeFA.
+BeFA outperformance than other parameter tuning methods.
 We visualize and analyze the content features adapted by LoRA and Soft-Prompt Tuning, as shown in Figure 6.
-The results indicate that while LoRA and Soft-Prompt Tuning provide some optimization, the degree of adaptation is limited.
-The adapted content features still exhibit issues such as content error and content insufficiency, failing to focus accurately on the content of the product itself.
-These methods fail to adequately address the issues of information drift and information omission.
-This is the reason for the overall poor performance of existing adaptation methods in recommendation tasks.
-In contrast, the impact of BeFA on content features content features is significantly better than that of existing methods, which is why BeFA performs well in the recommendation task.
-(d) Analysis in different scenarios
+The results indicate that although LoRA and Soft-Prompt Tuning offer some optimization, the degree of adaptation is limited.
+These methods do not sufficiently address issues of information drift and omission, failing to focus accurately on the content of the item itself.
+This explains why existing adaptation methods perform poorly in recommendation tasks.
+In contrast, BeFA has a greater impact on content characterisation than existing methods.
+The quality of its adjusted content features is obviously higher.
+It can better model the items and thus improve the performance.
+BeFA performs well across various scenarios.
 We employed three datasets ranging from simple to complex scenarios.
-The H&M dataset comprises only product images with minimal interference, making it relatively clean.
+The H&M dataset comprises only item images with minimal interference, making it relatively clean.
 In contrast, the Microlens dataset comprises more complex images with more background interference.
-Our adapter demonstrated significant improvements across both datasets.
-Specifically, on the H&M dataset, it increases $\mathrm{Recall@20}$ and $\mathrm{NDCG@20}$ by 14.17% and 17.34%, respectively.
-Meanwhile, on the Microlens dataset, our adapter increases $\mathrm{Recall@20}$ and $\mathrm{NDCG@20}$ by 4.58% and 5.19%, respectively.
-These findings indicate the adapter’s effectiveness across varying levels of scene complexity.
-Moreover, as illustrated in Figure 8, even in complex scenarios, our adapter effectively removes irrelevant information and locates the recommended products as precisely as possible
-5.3.2. Representation distribution analysis
-To visually demonstrate the impact of BeFA on adjusting content features, we visualize the distribution of visual representations in the TMALL dataset.
-Figure 9 illustrates the comparison of our adapter’s application to feature distributions across various recommendation models.
-Specifically, we randomly sample 500 item representations from the TMALL dataset and map them to two-dimensional normalized vectors on the unit hypersphere $\mathcal{S}^{1}$ (i.e., a circle with a radius of 1) by using t-SNE (tsne, ).
-Subsequently, we employ Gaussian kernel density estimation(KDE) (KDE, ) to plot the distribution of these two-dimensional features.
-Further examining the changes in the distribution of two dimensional features across different recommendation models before and after applying BeFA, we observe that the original feature distribution is more dispersed, forming multiple small dense regions.
-This distribution indicates that there is a large amount of noise in the visual feature extraction process, hindering the models’ ability to effectively distinguish key features of different products.
-In contrast, after applying BeFA, the feature distribution becomes more uniform and smooth, which indicates that our adapter has significant effects in eliminating noise.
-This improvement not only enhances feature recognition capability but also improves the model’s ability to distinguish different content features (pre, ).
-Moreover, comparing the changes in feature distributions across different models, it can be found that BeFA has strong adaptability.
-The feature distributions of various multimodal recommendation models show significant improvements after applying BeFA.
-This indicates that our adapter is not only effective on a single model but also has strong generalization capabilities across different recommender systems.
-This adaptability highlights BeFA’s wide potential and value in practical applications.
-5.4. Sensitivity Analysis(RQ4)
-Figure 10. Performance comparison w.r.t. different size of the decoupling space relative to the original embedding space $\lambda$.
-[…]
-We investigate the size of the decoupling space relative to the original embedding space, as shown in Fig.10.
+Our adapter demonstrated improvements across both datasets.
+Specifically, on the H&M dataset, it increases Recall@20 and NDCG@20 by 14.17% and 17.34%.
+Meanwhile, on the Microlens dataset, our adapter increases Recall@20 and NDCG@20 by 4.58% and 5.19%.
+BeFA reduces the interference between content features and filters and retains the most relevant features for the recommendation task, as illustrated in Figure 7.
+As a result, it achieves better recommendation results even in various complex situations.
+Hyperparameter Analysis
+Due to the length of the article, specific analyses are provided in the Appendix.
+Conclusion
+In this paper, we introduce an attribution analysis for visualization analyzing deficiencies of the content features.
+We found that the content features suffer from information drift and information omission, which lead to a decrease in the performance of the recommender systems.
+To address these issues, we propose Behavior-driven Feature Adapter (BeFA) which refine content features through the guidance of behavioral information.
+In our future work, we consider further designing more effective parameter-efficient adaptation methods.
+Appendix A Appendix
+Theoretical Analysis
+In this section, we reveal the deficiencies of content features extracted by pre-trained feature encoders within the embedding space.
+Furthermore, we demonstrate that optimizing the consistency between the extracted features and the ideal features necessary for accurate recommendation predictions can significantly enhance recommendation performance.
+This analysis provides theoretical support for the effectiveness of our proposed adapter.
+Analyzing features in the embedding space enables a better understanding and interpretation of their relationships.
+Formally, considering an item ${i}$ and its representation $\boldsymbol{e_{i,m}}$ in the embedding space, which represents the ideal content feature required by the recommendation model.
+On the basis of the item feature representation $\boldsymbol{e_{i,m}}$ in the embedding space, the recommender system calculates the probability distribution of the rating or clicking behavior of the user $u$ on the item $i$, which can be viewed as the prediction $P(r_{ui}|\boldsymbol{e_{i,m}},\boldsymbol{h_{i}},\boldsymbol{h_{u}})$.
+Where $r_{ui}$ stands for the rating or click behavior of the user $u$ on item $i$.
+Our objective is to find the optimal content feature $\boldsymbol{e_{i,m}}$ to maximize the posterior distribution.
+This can be formally described by the following objective function:
+$L(\boldsymbol{e_{i,m}})=\arg\max_{\boldsymbol{e_{i,m}}}P(r_{ui}|\boldsymbol{e_{i,m}},\boldsymbol{h_{i}},\boldsymbol{h_{u}})$ (8)
+The representation $\boldsymbol{e_{i,m}^{\prime}}$ extracted by the feature encoder can be seen as the prior distribution.
+However, there are deviations between the pre-extracted features and the ideal features.
+Let us denote this deviation as $\theta$.
+$\theta=\arccos\left(\frac{\boldsymbol{e_{i,m}}\cdot\boldsymbol{e_{i,m}^{\prime}}}{\|\boldsymbol{e_{i,m}}\|\|\boldsymbol{e_{i,m}^{\prime}}\|}\right)$ (9)
+A smaller $\theta$ indicates that the two representations are very close in the embedding space.
+This implies a high consistency between the pre-extracted features and the ideal features, which benefits the recommendation system in accurately capturing user interests and item similarities.
+By reducing $\theta$, the features become closer to the ideal features, thereby enhancing the performance of the recommendation system.
+The part $p$ of the representation $\boldsymbol{e_{i,m}^{\prime}}$ extracted by the encoder that is truly effective for the recommendation task can be represented as follows:
+$p=\boldsymbol{e_{i,m}^{\prime}}\cdot\left(\frac{\boldsymbol{e_{i,m}}\cdot\boldsymbol{e_{i,m}^{\prime}}}{|\boldsymbol{e_{i,m}}||\boldsymbol{e_{i,m}^{\prime}}|}\right)=\boldsymbol{e_{i,m}^{\prime}}\cdot\cos(\theta)$ (10)
+When the deviation $\theta$ is large, as $\boldsymbol{e_{i,m}^{\prime}}$ in Figure A1, content features exhibit the issues of information omission, causing the effective length of $p$ deviates significantly from $\boldsymbol{e_{i,m}}$.
+This results in a lack of crucial information in the extracted features, reflected as insufficient region of interest in the heatmap.
+Conversely, when $\theta$ is too large,as $\boldsymbol{e_{j,m}}$ in Figure A1, content features exhibit the issues of information drift, with $p$ being located in the wrong quadrant and its effective direction opposing $\boldsymbol{e_{i,m}}$.
+This introduces a large amount of incorrect information, reflected as incorrect region of interest in the heatmap.
+In such circumstances, the recommendation system is easily to make incorrect recommendations, because the feature representation contradicts the users’ preferences, failing to appropriately reflect the characteristics of items or users.
+In the recommendation task, we aim for a smaller deviation between $\boldsymbol{e_{i,m}^{\prime}}$ and $\boldsymbol{e_{i,m}}$.
+To formalize this objective, we define an error function $D(\boldsymbol{e_{i,m}^{\prime}},\boldsymbol{e_{i,m}})$ to measure the deviation between the representations:
+$D(\boldsymbol{e_{i,m}^{\prime}},\boldsymbol{e_{i,m}})=\left(1-\frac{{\boldsymbol{e_{i,m}^{\prime}}\cdot\boldsymbol{e_{i,m}}}}{{\|\boldsymbol{e_{i,m}^{\prime}}\|\|\boldsymbol{e_{i,m}}\|}}\right)$ (11)
+We seek to minimize the expected deviation measure $\Delta$. We can define $\Delta$ as:
+$\Delta=\mathbb{E}_{P(\boldsymbol{e_{i,m}^{\prime}})}[D(\boldsymbol{e_{i,m}^{\prime}},\boldsymbol{e_{i,m}})]$ (12)
+A decline in the quality of $\boldsymbol{e_{i,m}^{\prime}}$, indicated by an increase in the value of the error function $D(\boldsymbol{e_{i,m}^{\prime}},\boldsymbol{e_{i,m}})$, donates an augmented diversity between the representation extracted by the pre-trained encoder $\boldsymbol{e_{i,m}^{\prime}}$ and the ideal representation $\boldsymbol{e_{i,m}}$.
+Consequently, this leads to an increase in the expected deviation $\Delta$, subsequently influencing the posterior distribution $P(r_{ui}|\boldsymbol{e_{i,m}},\boldsymbol{h_{i}},\boldsymbol{h_{u}})$.
+Such circumstances may impair the recommendation system’s ability to accurately capturing the users’ preferences, thereby reducing the accuracy of recommendations.
+By minimizing $\Delta$, we can ensure that the representation $\boldsymbol{e_{i,m}^{\prime}}$ are closer to the ideal representation $\boldsymbol{e_{i,m}}$ in expectation.
+This adaptation of content features enhances the consistency with the ideal representation and thus improves the performance of the recommender system.
+Figure A1: An illustration of a theoretical analysis of deficiency analysis
+Experiments
+Datasets
+Dataset #User #Item #Behavior Density
+TMALL 13,104 7,848 151,928 0.148%
+Microlens 46,420 14,079 332,730 0.051%
+H&M 43,543 16,915 369,945 0.050%
+Table A1: Statistics of the experimental datasets
+We conducted experiments on three publicly available datasets: (a) TMALL [^4] https://tianchi.aliyun.com/dataset/140281; (b) Microlens [^5] https://recsys.westlake.edu.cn/MicroLens-50k-Dataset/; and (c) H&M [^6] https://www.kaggle.com/datasets/odins0n/handm-dataset-128x128.
+All these three datasets are real-world datasets.
+We performed 10-core filtering on the raw data.
+The detailed information of the filtered data is presented in Table A1.
+Evaluation Protocols
+For a fair comparison, we follow the evaluation settings in (Zhang et al.
+2021; Zhou et al.
+2023c) with the same 8:1:1 data splitting strategy for training, validation and testing.
+Besides, we follow the all-ranking protocol to evaluate the top-K recommendation performance and report the average metrics for all users in the test set: $\mathrm{R@K}$ and $\mathrm{N@K}$, which are abbreviations for $\mathrm{Recall@K}$ (Powers 2020) and $\mathrm{NDCG@k}$ (Järvelin and Kekäläinen 2002), respectively.
+Implementation
+We implement MMRec [^7] https://github.com/enoche/MMRec (Zhou 2023) based on PyTorch, which is a unified public repository designed for multimodal recommendation methods.
+To ensure fair comparison, we employed the Adam optimizer to optimize all methods and referred to the best hyperparameter settings reported in the original baseline paper.
+For general settings, we initialized embeddings with Xavier initialization of dimension 64, set the regularization coefficient to $\lambda_{E}=10^{-4}$, and batch size to $B=2048$.
+Early stopping and total epochs are fixed at 10 and 1000, respectively.
+We selecte the best model with the highest $\mathrm{Recall@20}$ metric on the validation set and reported metrics on the test set accordingly.
+Our experiments are done using RTX 4090 on windows system.
+Hyperparameter Analysis
+We investigate the size of the decoupling space relative to the original embedding space, as shown in Fig.A2.
 The results show that the optimal size of the decoupling space is about four times the size of the original embedding space on some datasets.
@@
 Additionally, its number of parameters and the computational cost will increase dramatically, bringing additional burden to the training process, which is impractical for practical application scenarios.
+Figure A2: Performance comparison w.r.t. different size of the decoupling space relative to the original embedding space $\lambda$.
 Furthermore, we find that there is a significant difference in the performance of Microlens, TMALL and H&M datasets under varying decoupling space sizes.
-For the Microlens dataset, both $\mathrm{Recall@20}$ and $\mathrm{NCGD@20}$ achieve the optimal performance when the decoupling space size is four times of the original embedding space.
+For the Microlens dataset,the optimal performance is achieved when the decoupling space size is four times of the original embedding space.
 In contrast, in the TMALL dataset, although the overall trend is similar, the optimal point fluctuates slightly, suggesting that the size of the decoupling space may need to be adjusted on different datasets to obtain optimal results.
@@
 This approach could adaptively adjust the decoupling space size for different application scenarios to achieve the best performance.
-6. Conclusion
-In this paper, we introduce an attribution analysis for visually analyzing deficiencies of the content features.
-We found that the content features of certain products suffer from information drift and information omission, and theoretically prove that these deficiencies lead to a decrease in the performance of the recommender systems.
-To address these issues, we propose Behavior-driven Feature Adapter (BeFA) adapter which adapts content features through the guidance of behavioral information, thereby enhancing the expressiveness of the content features and significantly improve the performance of the recommender systems
-In our future work, we consider introducing the size of the decoupling space as a learnable parameter into the training of the adapter, so that the optimized performance can be achieved in different scenarios, enhancing its general applicability.
+Appendix B Acknowledgment
+This work was supported the National Nature Science Foundation of China under Grants (No.62325206, 72074038), the Key Research and Development Program of Jiangsu Province under Grant BE2023016-4, the Natural Science Foundation of Jiangsu Province (BK.20210595) and the Postgraduate Research & Practice Innovation Program of Jiangsu Province under Grant (KYCX23_1026).